### 04 Mar 2026 v1

# Contribution of remote bands to orbital magnetization in twisted bilayer graphene

Abstract
Motivated by recent theoretical and experimental works on orbital magnetization $M_{\mathrm{orb}}$ for the interacting system, we develop a gauge-invariant framework to compute $M_{\mathrm{orb}}$ for correlated phases of magic-angle twisted bilayer graphene within self-consistent Hartree-Fock approximation. Based on the projector formulation of the theory of orbital magnetization, we evaluate both $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the Hartree–Fock Hamiltonian. We demonstrate that, in contrast to topological invariants such as the Chern number, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ obtain substantial contributions from remote bands and thus require careful convergence with respect to the number of included remote bands. Applying this approach to correlated phases at integer fillings, we obtain converged $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for time reversal symmetry broken Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at other integer fillings. Our results establish a systematic and controlled approach for evaluating orbital magnetization in correlated moiré systems and clarify the crucial role of remote bands in determining their magnetic response.
I Introduction
The recent discovery of the correlated insulating states Cao et al. (2018a); Sharpe et al. (2019); Lu et al. (2019); Jiang et al. (2019); Kerelsky et al. (2019); Xie et al. (2019); Choi et al. (2019); Serlin et al. (2020); Yankowitz et al. (2019) and unconventional superconductivity Cao et al. (2018b); Lu et al. (2019); Yankowitz et al. (2019); Codecido et al. (2019) in twisted bilayer graphene (TBG) has stimulated tremendous research efforts in moiré systems. Due to the extreme flatness of the lowest energy bands in TBG near the magic angle Bistritzer and MacDonald (2011) and their nontrivial topology Song et al. (2019, 2021); Ahn et al. (2019); Tarnopolsky et al. (2019); Liu et al. (2019a); Po et al. (2019), electronic Coulomb interactions are believed to drive a rich and robust set of strongly correlated and topological phases. Theoretical studies have predicted several strongly correlated and topologically distinct states at integer fillings  Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019). Among these states, the time-reversal symmetry breaking (TRSB) phases can become Chern insulators, exhibiting the quantum anomalous Hall (QAH) effect, which have been observed experimentally Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022). The ferromagnetic QAH states at $\nu=3$ filling was discovered to show strong magnetic response to external DC currents Sharpe et al. (2019); Serlin et al. (2020), which can be understood from the large orbital magnetization of the TBG system as a result of TRSB He et al. (2020); Su and Lin (2020); Huang et al. (2021). Specifically, this strong orbital effect has been studied in moiré systems where the time-reversal symmetry is explicitly broken Lee et al. (2019); Zhang et al. (2022), and in the $C_{2}\mathcal{T}$ symmetric moiré system such as TBG in the presence of additional $C_{2}\mathcal{T}$ symmetry breaking potential He et al. (2020); Su and Lin (2020).
Moreover, the transport and scanning tunneling microscopy (STM) experiments revealed that weak to moderate out-of-plane magnetic fields can turn the ground states at $\nu=\pm 1,\pm 2,\pm 3$ into strong-correlation induced Chern insulators in magic-angle TBG Das et al. (2021); Nuckolls et al. (2020). The associated orbital magnetization plays a central role in these phenomena, which determines the coupling to external magnetic fields. Understanding orbital magnetization in correlated moiré systems is therefore essential for elucidating the microscopic origin of Chern insulators and field-induced quantum phase transitions in TBG Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022).
Theoretically, the Berry curvature and orbital magnetic moments of moiré bands have been extensively analyzed within the non-interacting Bistritzer–MacDonald (BM) continuum model Liu et al. (2019b); He et al. (2020); Liu and Dai (2021b); Zhu et al. (2020). These studies established that the flat bands of TBG can host nontrivial topology and substantial orbital magnetization per unit cell largely arising from sizable Berry curvature. However, the flat-band nature of magic-angle TBG implies that electron–electron interactions are comparable to or even larger than the bandwidth, and self-consistent Hartree–Fock (HF) calculations have demonstrated that interactions tremendously reconstruct the band structure and stabilize correlated Chern insulating states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019). Despite this progress, a systematic calculation of the bulk orbital magnetization in such interaction-driven states remains lacking. In particular, it is not clear how interaction-induced band reconstruction, valley polarization, and remote-band contributions quantitatively modify the orbital magnetization from the non-interacting predictions, nor how these effects relate to experimentally observed field-induced switching between competing correlated phases.
In this work, we adopt the recently-developed method Kang et al. (2025); Zhu and Huang (2025); Liu et al. (2025); Ye (2026); Chen and Song (2026) and develop a gauge-invariant framework Ceresoli et al. (2006) for computing orbital magnetization within self-consistent Hartree–Fock theory and apply it to correlated insulating states of magic-angle TBG. Starting from the BM continuum model with Coulomb interactions projected onto the relevant bands, we determine symmetry-broken ground states at integer fillings and evaluate the orbital magnetization using the modern theory expressed in terms of projection operators. This formulation avoids gauge ambiguities and enables a controlled treatment of the large number of remote bands inherent to the continuum description. We demonstrate that, unlike topological invariants such as the Chern number, the orbital magnetization exhibits a pronounced sensitivity to remote-band contributions and requires careful convergence. Applying this framework, we compute both the total orbital magnetization $M_{orb}$ and the self-rotation (local circulation) component $m_{SR}$ relevant for magneto-optical measurements such as magnetic circular dichroism (MCD). The calculated orbital magnetization respects the particle-hole symmetry that relates the fillings of $\nu$ and $\nu$ Bernevig et al. (2021) and reproduces the characteristic linear dependence of magnetization on chemical potential within the insulating gap, with slope proportional to the Chern number. We further analyze the sign structure of the magnetization and its relation to changes in Hall conductivity. At $\nu=\pm 1$, we compare competing Chern and intervalley-coherent states and show that Chern insulating phases possess substantially larger orbital magnetization, providing a microscopic explanation for their stabilization under external magnetic fields. Our results establish a systematic approach for evaluating orbital magnetization in correlated moiré systems and elucidate the important role of interactions as well as the critical contribution of remote bands in shaping orbital ferromagnetism in twisted bilayer graphene.
II Non-interacting model
The moiré pattern in TBG is formed by twisting the top and bottom layers of an aligned bilayer graphene by angles $\theta/2$ and $\theta/2$, which forms a periodic superlattice structure of a triangular lattice with the lattice vectors $\mathbf{L_{M1}}=(0,-1)L_{M}$, $\mathbf{L_{M2}}=(\frac{\sqrt{3}}{2},-\frac{1}{2})L_{M}$ and $L_{M}=a_{0}/(2\sin(\theta/2))$, where $a_{0}$=0.246 nm, is the lattice constant of the monolayer graphene. The corresponding reciprocal lattice vectors of the moiré lattice are ${\mathbf{G}}_{1}=(-\frac{2\pi}{\sqrt{3L_{M}}},-\frac{2\pi}{L_{M}})$ and ${\mathbf{G}}_{2}=(\frac{4\pi}{\sqrt{3L_{M}}},0)$. The non-interacting physics of TBG can be described by the continuum model introduced by Bistritzer and MacDonald Bistritzer and MacDonald (2011); Lopes dos Santos et al. (2012). The Hamiltonian of the continuum model for each valley is given by,
$H^{\tau}_{BM}(\hat{\mathbf{k}})=\left(\begin{array}[]{cc}-\hbar v_{F}(\hat{{\mathbf{k}}}-{\mathbf{K}}_{1}^{\tau})\cdot\boldsymbol{\sigma}^{\tau}&U\\ U^{+}&-\hbar v_{F}(\hat{{\mathbf{k}}}-{\mathbf{K}}_{2}^{\tau})\cdot\boldsymbol{\sigma}^{\tau}\end{array}\right)$ (1)
where $\hat{{\mathbf{k}}}=-i\partial_{{\mathbf{r}}}$, $\tau=\pm$ is the valley index, and $\boldsymbol{\sigma}^{\tau}=(\tau\sigma_{x},\sigma_{y})$ the Pauli matrices acting in the A,B sublattice space. In Eq. (1), ${\mathbf{K}}^{\tau}_{1}$ and ${\mathbf{K}}^{\tau}_{2}$ are the Dirac points of valley $\tau$ in the bottom and top layers that are twisted by angles $\mp\frac{\theta}{2}$, which are equivalent to $\tau{\mathbf{K}}_{1}$ and $\tau{\mathbf{K}}_{2}$ in the mBZ as shown in Fig. 1(c). The interlayer tunneling between the Dirac states in the two layers is described by the matrix
$\begin{split}U=&\left(\begin{array}[]{cc}u_{0}&u_{1}\\ u_{1}&u_{0}\end{array}\right)+\left(\begin{array}[]{cc}u_{0}&u_{1}e^{-i\tau\frac{2\pi}{3}}\\ u_{1}e^{i\tau\frac{2\pi}{3}}&u_{0}\end{array}\right)e^{-i\tau{\mathbf{G}}_{1}\cdot{\mathbf{r}}}\\ &+\left(\begin{array}[]{cc}u_{0}&u_{1}e^{i\tau\frac{2\pi}{3}}\\ u_{1}e^{-i\tau\frac{2\pi}{3}}&u_{0}\end{array}\right)e^{-i\tau({\mathbf{G}}_{1}+{\mathbf{G}}_{2})\cdot{\mathbf{r}}}\end{split}$ (2)
where $u_{0}$ and $u_{1}$ are the intra- and inter-sublattice interlayer tunneling amplitudes. We adopt realistic parameters for the continuum model: $\hbar v_{F}/a_{0}=2.365$ eV, inter-sublattice interlayer coupling $u_{1}=0.11$ eV, and intra-sublattice interlayer coupling $u_{0}=0.06$ eV. The twist angle is fixed at the magic value $\theta$=1.086, at which the low-energy moiré bands become perfectly flat in the limit $u_{0}\to 0$ Tarnopolsky et al. (2019). The low-energy flat bands exhibit a bandwidth of approximately 2 meV and are well separated from the higher-energy remote bands by an energy gap of about 50 meV as shown in Fig. 1. The eigenstates of the continuum Hamiltonian $H_{BM}^{\tau}$ have the Bloch wavefucntion form
$\psi_{m,\tau,{\mathbf{k}}}^{X}({\mathbf{r}})=\sum_{{\mathbf{G}}}u_{m,\tau;{\mathbf{G}},X}({\mathbf{k}})e^{i({\mathbf{k}}+{\mathbf{G}})\cdot{\mathbf{r}}}$ (3)
where $X=\{A_{1},B_{1},A_{2},B_{2}\}$ is the layer and sublattice index with the eigen-band energy $\epsilon_{m{\mathbf{k}}\tau}$ and $\mathbf{G}$} is the general reciprocal lattice vector in the mBZ. Here, $m$ and $\tau$ are the band and valley indices, and the spin index $s$ is omitted since the Hamiltonian is spin independent.
Figure 1: (a) Low-energy band structure of non-interacting TBG calculated using the BM model. The black solid and red dashed curves correspond to the $K$’ and $K$ valleys, respectively. (b) Zoomed-in view of the isolated flat bands near charge neutrality, highlighting their separation from the higher-energy remote bands. (c) Schematic illustration of Brillouin zone folding in TBG. The small hexagon denotes the moiré Brillouin zone (mBZ), reciprocal to the moiré superlattice. The cyan arrows indicate the momentum path in the enlarged moiré Brillouin zone along which the band dispersions in panels (a) and (b) are plotted.
III Interacting case
The full Hamiltonian $H=H_{0}+H_{int}$ contains the non-interacting part
$H_{0}=\sum_{\tau}H_{BM}^{\tau}=\sum_{m{\mathbf{k}}s\tau}\epsilon_{m{\mathbf{k}}\tau}d_{m{\mathbf{k}}s\tau}^{\dagger}d_{m{\mathbf{k}}s\tau},$ (4)
where $d^{\dagger}_{m{\mathbf{k}}s\tau}$ and $d_{m{\mathbf{k}}s\tau}$ are the creation and annihilation operators in the eigen-band basis. To incorporate interactions, we project the Coulomb interaction onto the two lowest-energy flat bands per spin and per valley, and perform self-consistent Hartree–Fock calculations directly in momentum space. This approach avoids the real-space Wannier obstruction associated with topological flat bands Po et al. (2019, 2018a, 2018b); Ahn et al. (2019); Song et al. (2021). Further technical details of the Hartree–Fock procedure can be found in the Appendix of Ref.Zhang et al., 2020. The interacting part of the Hamiltonian is written asBernevig et al. (2021); Kang et al. (2021)
$H_{int}=\frac{1}{2S}\sum_{{\mathbf{q}}\neq 0}V({\mathbf{q}})\delta\rho_{{\mathbf{q}}}\delta\rho_{-{\mathbf{q}}}\ ,$ (5)
where S is the total area and $\delta\rho_{{\mathbf{q}}}=\rho_{{\mathbf{q}}}-\bar{\rho}_{{\mathbf{q}}}$ is the density operator relative to charge neutrality of the decoupled bilayer graphene in momentum space. We note that this interaction form is somewhat different from the normal ordered interactions supplemented by double-counting subtraction schemes used in previous Hartree-Fock (HF) studies Bultinck et al. (2020b); Liu et al. (2021); Zhang et al. (2020). It is believed that the Coulomb interaction expressed in Eq. 5 has the advantage of partially accounting for the renormalization and screening effects of the remote bands on the lowest flat bands Kang et al. (2021); Bernevig et al. (2021). The density operator can be expanded in the eigen-basis of $H_{0}$,
$\rho_{{\mathbf{q}}}=\sum_{s,\tau}\sum_{{\mathbf{k}}\in\text{mBZ},m_{1},m_{2}}\lambda_{m_{1},m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})d^{\dagger}_{{\mathbf{k}},m_{1},s,\tau}d_{{\mathbf{k}}+{\mathbf{q}},m_{2},s,\tau}\,$ (6)
with the quantum averaged background density
$\bar{\rho}_{{\mathbf{q}}}=\sum_{{\mathbf{G}},\tau,m}\delta_{{\mathbf{q}},{\mathbf{G}}}\sum_{{\mathbf{k}}\in\text{mBZ}}\lambda_{m,m;\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{G}})\ ,$ (7)
where $\lambda$ is the form factor defined in terms of the Bloch wavefunction as
$\begin{split}\lambda_{m_{1}m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})&=\left<u_{m_{1},\tau}({\mathbf{k}})|u_{m_{2},\tau}({\mathbf{k}}+{\mathbf{q}})\right>\\ &=\sum_{{\mathbf{G}}^{{}^{\prime}},X}u^{*}_{m_{1},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}})u_{m_{2},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}}+{\mathbf{q}}).\end{split}$ (8)
Here, $\mathbf{k}$} is restricted in the mBZ while $\mathbf{q}$} can extend to the original Brillouin zone of the monolayer graphene. The operator $d^{\dagger}$ obeys the periodic condition $d^{\dagger}_{m{\mathbf{k}}s\tau}=d^{\dagger}_{m{\mathbf{k}}+{\mathbf{G}}s\tau}$, which is equivalent to the Bloch wave function obeying the condition $u_{m,\tau;{\mathbf{G}},X}({\mathbf{k}}+{\mathbf{G}}_{0})=u_{m,\tau;{\mathbf{G}}+{\mathbf{G}}_{0},X}({\mathbf{k}})$. We take $V({\mathbf{q}})$ as the single-gate-screened Coulomb potential Bultinck et al. (2020b)
$V({\mathbf{q}})=\frac{e^{2}}{2\varepsilon\varepsilon_{0}q}(1-e^{-2qd_{s}})$ (9)
In this paper, we take the dielectric constant $\varepsilon=7$ and the gate distance $d_{s}=40$ nm.
It is well established that electron–electron Coulomb interactions in twisted bilayer graphene can drive spontaneous symmetry breaking, giving rise to a series of correlated insulating states at commensurate integer fillings. Among these, states that break the combined $C_{2}\mathcal{T}$ symmetry are energetically competitive. In particular, at fillings $\nu=\pm 3$, the $C_{2}\mathcal{T}$-broken Chern insulating states can emerge as ground states even in the absence of an explicit staggered sublattice potential. In this work, we systematically identify the correlated insulating states at all integer fillings within the Hartree–Fock approximation and compute the orbital magnetization of these interaction-driven phases.
IV Orbital magnetization
The topology of the correlated electronic states and their response to an external magnetic field are governed by the Berry curvature and orbital magnetization of the interaction-renormalized bands. In our approach, the Coulomb interaction is projected onto the lowest two flat bands per spin and per valley. As a consequence, only those HF renormalized flat bands that break the combined $C_{2}\mathcal{T}$ symmetry can acquire nontrivial Berry curvature and hence finite Chern numbers. Since the Berry curvature depends only on the periodic part of the Bloch wave functions, it can be directly computed from the HF eigenstates as
${\bf\Omega}_{n}({\mathbf{k}})=i\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}^{HF}(\bm{k})\middle|\partial_{\nu}u_{n}^{HF}(\bm{k})\right\rangle,$ (10)
where $n$ denotes the band index, $\epsilon_{\mu\nu}$ is the antisymmetric tensor, $\mu,\nu$ are Cartesian indices, and $\partial_{\mu}\equiv\partial/\partial k_{\mu}$. For the orbital magnetization, it has been recently shown Kang et al. (2025) that both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ can be evaluated using the HF Hamiltonian and its eigenfunctions Souza and Vanderbilt (2008); Resta (2020); Ji et al. (2025):
$M_{orb}=\frac{e}{2i\hbar}\sum_{n\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\left|\left(H^{HF}(\bm{k})+\epsilon_{n}^{HF}(\bm{k})-2\mu\right)\right|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}(\epsilon_{n}^{HF}(\bm{k})-\mu)$ (11)
$m_{SR}=\frac{e}{2i\hbar}\sum_{n,\alpha}\sum_{\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\middle|u_{\alpha}(\bm{k})\right\rangle\left(\epsilon^{HF}_{\alpha}(\bm{k})-\epsilon_{n}^{HF}(\bm{k})\right)\left\langle u_{\alpha}(\bm{k})\middle|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(\epsilon_{n}^{HF}(\bm{k})-\mu\right)\left(1-n_{F}(\epsilon_{\alpha}^{HF}(\bm{k})-\mu)\right)$ (12)
where $n_{F}(E)=\Theta(\mu-E)$ is the Fermi–Dirac distribution function at $T=0$ and $\mu$ is the chemical potential. These quantities can be reformulated in terms of projection operators.
We emphasize that the Hartree-Fock Hamiltonian $H^{\mathrm{HF}}(\mathbf{k})$ must be formulated in the full Hilbert space, including the flat-band sector as well as both the occupied and empty remote bands. In principle, Eqs. 11 and  12 involve summations over all occupied states in both the active and remote bands, rendering the calculation formally infinite. To make the numerical calculations tractable, we evaluate $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ in a truncated Hilbert space. For clarity of notation, we label the HF Bloch states $|u_{n}^{A/R+/R-}(\mathbf{k})\rangle$ by the superscripts $A$, $R+$, and $R$-, which denote the active bands, the empty remote bands, and the occupied remote bands, respectively. The empty and occupied remote bands are indexed by $n$, counted outward from the charge neutrality point, such that the states $|u_{n}^{R+}(\mathbf{k})\rangle$ and $|u_{n}^{R-}(-\mathbf{k})\rangle$ are related by the particle-hole symmetry Bernevig et al. (2021). On this basis, we introduce two projection operators onto the truncated occupied and empty states.
$P_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{occupied}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R-}_{n}(\mathbf{k})\rangle\langle u^{R-}_{n}(\mathbf{k})|$ (13)
$Q_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{empty}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R+}_{n}(\mathbf{k})\rangle\langle u^{R+}_{n}(\mathbf{k})|,$ (14)
These two projection operators for truncated occupied and empty states explicitly depend on the choice of $n_{\rm cut}$, and thus $P_{n_{\rm cut}}(k)+Q_{n_{\rm cut}}(k)=I$ only when $n_{\rm cut}\rightarrow\infty$. In this representation, we introduce the orbital magnetization for the truncated Hilbert space that is expressed as sm
$M^{\rm trun}_{orb}=-\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W^{\rm trun}_{xy}(\bm{k})-N^{\rm trun}_{xy}(\bm{k})\right]$ (15)
$m^{\rm trun}_{SR}=\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W^{\rm trun}_{xy}(\bm{k})+N^{\rm trun}_{xy}(\bm{k})\right],$ (16)
where
$W^{\rm trun}_{\mu\nu}(\bm{k})=\mathrm{Tr}\left[P_{n_{cut}}(\bm{k})\big(\partial_{\mu}P_{n_{cut}}(\bm{k})\big)\big(\partial_{\nu}Q_{n_{cut}}(\bm{k})\big)\big(H^{HF}(\bm{k})-\mu\big)\right]$ (17)
and
$N^{\rm trun}_{\mu\nu}(\bm{k})=\mathrm{Tr}\left[Q_{n_{cut}}(\bm{k})\big(\partial_{\mu}P_{n_{cut}}(\bm{k})\big)\big(\partial_{\nu}Q_{n_{cut}}(\bm{k})\big)\big(H^{HF}(\bm{k})-\mu\big)\right].$ (18)
Because the projection operators $P_{n_{cut}}(\mathbf{k})$ and $Q_{n_{cut}}(\mathbf{k})$ are manifestly gauge invariant, this formulation avoids the gauge ambiguities that typically arise in direct evaluations of orbital magnetization based on Bloch wave functions. Moreover, it circumvents the singularities associated with band degeneracy points. Formally, physical quantities are defined in the limit $n_{\mathrm{cut}}\to\infty$. Nevertheless, we anticipate rapid convergence with increasing $n_{\mathrm{cut}}$, enabling reliable numerical evaluation within a finite truncated Hilbert space.
IV.1 $\nu=\pm 3$
We first consider the case of fillings $\nu=\pm 3$. The self-consistent Hartree–Fock calculation yields a spin–valley polarized Chern insulating ground state characterized by spontaneous breaking of the combined $C_{2}\mathcal{T}$ symmetry. The corresponding band dispersions, shown in Fig. 2(a) and (b), indicate that among the eight active bands, a single pair associated with one spin–valley flavor develops a gap due to $C_{2}\mathcal{T}$ symmetry breaking and acquires a finite Chern number. The remaining three pairs of bands, corresponding to the other spin–valley flavors (shown in black), remain topological trivial. The symmetry breaking generates an indirect energy gap of approximately 13 meV between the occupied Chern band and the higher-energy bands. These states correspond to the experimentally observed quantum anomalous Hall phases with Chern number $C=1$ (C=1C=-1) at fillings $\nu=3$ ($\nu=-3$Sharpe et al. (2019); Serlin et al. (2020); Nuckolls et al. (2020).
Figure 2: (a,b) Band dispersions of the Chern insulating states at fillings $\nu=3$ (a) and $\nu=-3$ (b), obtained within the HF approximation. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands that are excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1), corresponding to the experimentally observed Chern insulating states with $C=\pm 1$ at $\nu=\pm 3$. (c,d) Calculated total orbital magnetization $M_{\mathrm{orb}}$ and self-rotation contribution $m_{\mathrm{SR}}$ per moiré cell for the states shown in (a,b), plotted as functions of the number of remote-band pairs ($n_{\mathrm{cut}}$) included in the projection-matrix formalism. Both quantities exhibit clear convergence for $n_{\mathrm{cut}}\approx 20$. When evaluating $M_{\mathrm{orb}}$, the chemical potential $\mu$ is fixed at the top of the valence band for $\nu=3$ and at the bottom of the conduction band for $\nu=-3$, as indicated by the horizontal dashed lines in panels (a) and (b).
We next compute the orbital magnetization of these states using Eqs. 15 and 16. As shown in Fig.2(c,d), both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation component $m_{\mathrm{SR}}$ converge for $n_{\mathrm{cut}}\approx 20$. Moreover, the results respect the particle–hole symmetry relating fillings $\nu=3$ and $\nu=-3$, satisfying $M_{\mathrm{orb}}(\nu=3)=M_{\mathrm{orb}}(\nu=-3)$, and $m_{\mathrm{SR}}(\nu=3)=-m_{\mathrm{SR}}(\nu=-3)$ sm , consistent with the experimentally observed Landau fan structureDas et al. (2021); Nuckolls et al. (2020). Since convergence is achieved for $n_{\mathrm{cut}}\approx 20$, we adopt $n_{\mathrm{cut}}=30$ in the following calculations to ensure numerical stability.
One might be tempted to conclude that both the orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ arise exclusively from the occupied Chern bands that explicitly break time-reversal symmetry, and that only these bands need to be considered—analogous to the case of Berry curvature. Such reasoning, however, is not generally valid. Although the occupied states of the remote bands, $|u_{n}(\mathbf{k})\rangle$, remain invariant under time-reversal symmetry, their momentum derivatives $|\partial_{\mu}u_{n}(\mathbf{k})\rangle$ generally possess non-vanishing overlaps with those states in the active bands. Consequently, once time-reversal symmetry is broken within the active bands, the nominally time-reversal-symmetric remote bands acquire finite contributions to both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$. The need to include a substantial number of remote bands—at least $20$ pairs in our calculations—to achieve numerical convergence provides direct evidence of their non-negligible contributions.
More physically relevant is the behavior of $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ within the Chern insulating gap, indicated by the red dots in Fig. 3(a,b). The total orbital magnetization $M_{\mathrm{orb}}$ exhibits the expected linear dependence on $\mu$. Thermodynamics gives the relation $\partial M_{orb}/\partial\mu=\partial N/\partial B$, consistent with the Diophantine equation that characterizes quantum Hall states. Taking the derivative of Eq.11 with respect to $\mu$ yields
$\frac{\partial M_{orb}}{\partial\mu}=S\frac{e}{h}C,$ (19)
where $C$ is the Chern number. The linear dependence leads to a sign reversal of $M_{\mathrm{orb}}$ across the gap, corresponding to magnetization reversal previously analyzed in the non-interacting case Zhu et al. (2020). The in-gap plateau of $m_{\mathrm{SR}}$ demonstrates that the Chern insulating state possesses a large self-rotation contribution of approximately $40\mu_{B}$, which is directly accessible through magneto-optical probes such as magnetic circular dichroism.
Figure 3: Calculated total orbital magnetization $M_{\mathrm{orb}}$ (a) and self-rotation contribution $m_{\mathrm{SR}}$ (b) per moiré cell as functions of the chemical potential for the Chern insulating state at $\nu=3$ inside the insulating gap, whose band dispersion is shown in Fig. 2(a). $M_{\mathrm{orb}}$ shows a linear dependence with $\mu$ and changes sign within the gap while $m_{\mathrm{SR}}$ remains constant. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.
IV.2 Competing states at $\nu=\pm 1$
We next consider the correlated insulating states at fillings $\nu=\pm 1$. Owing to the particle–hole symmetry relating $\nu$ and $\nu$, we focus on the case $\nu=1$. The Hartree–Fock calculation yields two nearly degenerate correlated insulating states. The lowest-energy state is a mixed phase combining intervalley coherence (IVC) and $C_{2}\mathcal{T}$ symmetry breaking. In this state, the five valence bands consist of two bands associated with a spin-polarized IVC sector carrying zero Chern number, one $C_{2}\mathcal{T}$-broken band with finite Chern number, and an additional pair of ungapped bands from one spin–valley flavor, resulting in a total Chern number $C=1$ Zhang et al. (2020). Its band dispersion is shown in Fig.4(b). This state is lower in energy by approximately 0.5 meV per particle compared to the state in which three valence bands break $C_{2}\mathcal{T}$ symmetry and each carries a finite Chern number. We refer to the latter as the correlated Chern insulator (CCI) with total Chern number $C=3$, whose dispersion is shown in Fig.4(a). We then compute the orbital magnetization for both states. As shown in Fig.4(c), the total orbital magnetization $M_{\mathrm{orb}}$ for both states exhibits a linear dependence on the chemical potential within the insulating gap and changes sign across the gap. Both the magnitude and the slope of $M_{\mathrm{orb}}$ are larger for the CCI than for the mixed IVC state, reflecting its higher Chern number. This implies that the CCI is energetically favored in the presence of a finite magnetic field, consistent with experimental observations. Similarly, the self-rotation contribution $m_{\mathrm{SR}}$ of the CCI is significantly larger than that of the IVC state, as shown in Fig.4(d).
Figure 4: (a,b) Band dispersions of the correlated Chern insulating state (CCI) (a) and the intervalley-coherent (IVC) state (b) at filling $\nu=1$, obtained within the HF approximation. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). For the CCI, the Chern bands are triply degenerate, yielding a total Chern number $C=3$. (c,d) Calculated total orbital magnetization $M_{\mathrm{orb}}$ (c) and self-rotation contribution $m_{\mathrm{SR}}$ (d) per moiré cell as functions of the chemical potential $\mu$ inside the insulating gap for the two states shown in (a) and (b). Solid circles correspond to the CCI, and solid squares to the IVC state. $M_{\mathrm{orb}}$ varies linearly with $\mu$ and changes sign across the gap, while $m_{\mathrm{SR}}$ remains constant. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.
V Conclusion
In this work, we developed a practical and gauge-invariant framework to compute orbital magnetization in correlated states of magic-angle twisted bilayer graphene within self-consistent Hartree–Fock theory. Based on the projector formulation of the recently developed theory of orbital magnetization in the interacting system Kang et al. (2025), we evaluated both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the HF Hamiltonian and its eigenstates, thereby avoiding gauge ambiguities associated with Bloch-wavefunction derivatives. A key advantage of this approach is that it enables controlled inclusion of remote-band effects through the projection matrices. We demonstrated that, unlike the Chern number, the orbital magnetization is highly sensitive to remote bands and requires careful convergence with respect to the number of remote band pairs $n_{\mathrm{cut}}$, which saturates for $n_{\mathrm{cut}}\gtrsim 20$ in our calculations.
Applying this method to integer fillings, we obtained converged orbital magnetization for the TRSB Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at $\nu=1$. Our results ensure the thermodynamic linear dependence of $M_{\mathrm{orb}}$ on the chemical potential inside the insulating gaps with the slope determined by the Chern number, while yielding a sizable and constant in-gap $m_{\mathrm{SR}}$ relevant for magneto-optical probes. More broadly, the gauge-invariant projector formalism introduced here provides a systematic route to quantify orbital magnetization in interaction-driven moiré phases and can be readily extended to other moiré materials.
ACKNOWLEDGMENTS
We thank Hui-Ke Jin, Jie Wang, Minxuan Wang, Wen-Yu He, and especially Oskar Vafek for helpful discussions. J. K. acknowledges the support from the NSFC (Grant No. 12574159 and No. 12074276), the Double First-Class Initiative Fund of ShanghaiTech University, and the start-up grant of ShanghaiTech University. YZ is supported in part by the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800) and National Natural Science Foundation of China (NSFC) Grants No. 12274279. ZW is supported by the U.S. Department of Energy, Basic Energy Sciences, Grant No. DE FG02-99ER45747.