### 16 Apr 2026 v2

# Contribution of remote bands to orbital magnetization in twisted bilayer graphene

Abstract
Motivated by recent theoretical and experimental studies of orbital magnetization $M_{\mathrm{orb}}$ in interacting systems, we develop a gauge-invariant framework to compute $M_{\mathrm{orb}}$ for correlated phases of magic-angle twisted bilayer graphene within a self-consistent Hartree–Fock approximation. Based on the projector formulation of the theory of orbital magnetization, we evaluate both $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the Hartree–Fock Hamiltonian. We demonstrate that, in contrast to topological invariants such as the Chern number, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ receive substantial contributions from remote bands and thus require careful convergence with respect to the number of included remote bands. Applying this approach to correlated phases at integer fillings, we obtain converged $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for time-reversal-symmetry-broken Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at other integer fillings. Our results establish a systematic and controlled approach for evaluating orbital magnetization in correlated moiré systems and clarify the crucial role of remote bands in determining their magnetic response.
I Introduction
The discovery of correlated insulating states Cao et al. (2018a); Jiang et al. (2019); Kerelsky et al. (2019); Xie et al. (2019); Choi et al. (2019); Xie et al. (2021a); Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022); Lu et al. (2019); Yankowitz et al. (2019) and unconventional superconductivity Cao et al. (2018b); Lu et al. (2019); Yankowitz et al. (2019); Codecido et al. (2019) in twisted bilayer graphene (TBG) has stimulated intense interest in correlated moiré systems. Near the magic angle, the isolated low-energy bands are extremely flat Bistritzer and MacDonald (2011) and possess nontrivial topology Tarnopolsky et al. (2019); Liu et al. (2019a); Song et al. (2019, 2021); Ahn et al. (2019); Po et al. (2019, 2018a, 2018b), so electron-electron interactions are expected to drive a rich variety of symmetry-broken and topological phases. Indeed, theoretical studies have predicted several strongly correlated states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Lian et al. (2021); Xie et al. (2021b); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019); Bernevig et al. (2021); Kang et al. (2021); Zhang et al. (2020); Wang and Vafek (2024, 2022). Among these phases, time-reversal symmetry breaking (TRSB) states are particularly interesting because they can become Chern insulators and exhibit integer or fractional quantum anomalous Hall (QAH) effects, as observed experimentally in TBG Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022); Xie et al. (2021a) and other moiré materials Chen et al. (2020); Li et al. (2021); Xu et al. (2023); Park et al. (2023); Lu et al. (2024). In particular, the ferromagnetic QAH state at $\nu=3$ was found to exhibit a strong magnetic response to external dc currents Sharpe et al. (2019); Serlin et al. (2020), pointing to a large orbital magnetization associated with TRSB He et al. (2020); Su and Lin (2020); Huang et al. (2021); Lee et al. (2019); Zhang et al. (2022). Orbital magnetization in TBG has also been probed directly by high-resolution magnetometry Tschirhart et al. (2021); Grover et al. (2022).
Transport and scanning tunneling microscopy experiments further showed that weak to moderate out-of-plane magnetic fields can drive the states at $\nu=\pm 1,\pm 2,\pm 3$ into correlation-induced Chern insulators in magic-angle TBG Das et al. (2021); Nuckolls et al. (2020). These observations highlight orbital magnetization as a key quantity governing the coupling of correlated moiré states to external magnetic fields. A quantitative understanding of orbital magnetization is therefore essential for elucidating the microscopic origin of Chern-insulating phases and field-induced transitions between competing states in TBG Das et al. (2021); Nuckolls et al. (2020); Stepanov et al. (2021); Yu et al. (2022).
On the theoretical side, the Berry curvature and orbital magnetic moments of moiré bands have been extensively analyzed within the non-interacting Bistritzer–MacDonald (BM) continuum model Liu et al. (2019b); He et al. (2020); Liu and Dai (2021b); Zhu et al. (2020). These studies established that the flat bands of TBG can host nontrivial topology and substantial orbital magnetization per unit cell largely arising from sizable Berry curvature. However, the flat-band nature of magic-angle TBG implies that electron–electron interactions are comparable to or even larger than the bandwidth, and self-consistent Hartree–Fock (HF) calculations have demonstrated that interactions tremendously reconstruct the band structure and stabilize correlated Chern insulating states at integer fillings Kang and Vafek (2019, 2020); Zhang et al. (2019, 2020); Xu et al. (2018); Xie and MacDonald (2020); Liu et al. (2018); Huang et al. (2019); Liu and Dai (2021a); Liu et al. (2021); Bultinck et al. (2020a, b); Wu and Das Sarma (2020); Chen et al. (2021); Lu et al. (2022); Da Liao et al. (2021); Bernevig et al. (2021); Lian et al. (2021); Xie et al. (2021b); Hejazi et al. (2021); Sboychakov et al. (2019); Rademaker et al. (2019); Guinea and Walet (2018); Seo et al. (2019); Kang et al. (2021); Wang and Vafek (2024, 2022). Despite this progress, a systematic calculation of the bulk orbital magnetization in such interaction-driven states remains lacking. In particular, it is not clear how interaction-induced band reconstruction, valley polarization, and remote-band contributions quantitatively modify the orbital magnetization beyond the non-interacting predictions, nor how these effects relate to experimentally observed field-induced switching between competing correlated phases.
In this work, we combine the recently developed projector formulation of orbital magnetization Kang et al. (2025); Zhu and Huang (2025); Liu et al. (2025); Ye (2026); Chen and Song (2026) with self-consistent Hartree-Fock theory to construct a gauge-invariant framework for correlated states of magic-angle TBG Ceresoli et al. (2006). Starting from the BM continuum model with Coulomb interactions projected onto the relevant bands, we determine symmetry-broken ground states at integer fillings and evaluate the orbital magnetization using the modern theory expressed in terms of projection operators. This formulation avoids gauge ambiguities and enables a controlled treatment of the large number of remote bands inherent to the continuum description.
Using this framework, we compute both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ relevant for magneto-optical probes such as magnetic circular dichroism (MCD) Xia et al. (2026). We show that, unlike topological invariants such as the Chern number, the orbital magnetization is strongly sensitive to remote-band contributions and therefore requires careful convergence. The calculated orbital magnetization respects the particle-hole symmetry that relates the fillings of $\nu$ and $\nu$ Bernevig et al. (2021) and reproduces the characteristic linear dependence of magnetization on chemical potential within the insulating gap, with slope proportional to the Chern number. We further analyze the sign structure of the magnetization and its relation to changes in Hall conductivity. At $\nu=\pm 1$, we compare competing Chern and intervalley-coherent states and show that Chern insulating phases possess substantially larger orbital magnetization, providing a microscopic explanation for their stabilization under external magnetic fields. Our results establish a systematic approach for evaluating orbital magnetization in correlated moiré systems and elucidate the important role of interactions as well as the critical contribution of remote bands in shaping orbital ferromagnetism in twisted bilayer graphene.
II Non-interacting model
The moiré pattern in TBG is formed by twisting the top and bottom layers of an aligned bilayer graphene by angles $\theta/2$ and $\theta/2$, which forms a periodic superlattice structure of a triangular lattice with the lattice vectors $\mathbf{L_{M1}}=(0,-1)L_{M}$, $\mathbf{L_{M2}}=(\frac{\sqrt{3}}{2},-\frac{1}{2})L_{M}$ and $L_{M}=a_{0}/(2\sin(\theta/2))$, where $a_{0}$=0.246 nm, is the lattice constant of the monolayer graphene. The corresponding reciprocal lattice vectors of the moiré lattice are ${\mathbf{G}}_{1}=(-\frac{2\pi}{\sqrt{3}L_{M}},-\frac{2\pi}{L_{M}})$ and ${\mathbf{G}}_{2}=(\frac{4\pi}{\sqrt{3}L_{M}},0)$. The non-interacting physics of TBG can be described by the continuum model introduced by Bistritzer and MacDonald Bistritzer and MacDonald (2011); Lopes dos Santos et al. (2012). The Hamiltonian of the continuum model for each valley is given by,
$H^{\tau}_{BM}(\hat{\mathbf{k}})=\left(\begin{array}[]{cc}-\hbar v_{F}(\hat{{\mathbf{k}}}-{\mathbf{K}}_{1}^{\tau})\cdot\boldsymbol{\sigma}^{\tau}&U\\ U^{+}&-\hbar v_{F}(\hat{{\mathbf{k}}}-{\mathbf{K}}_{2}^{\tau})\cdot\boldsymbol{\sigma}^{\tau}\end{array}\right)$ (1)
where $\hat{{\mathbf{k}}}=-i\partial_{{\mathbf{r}}}$, $\tau=\pm$ is the valley index, and $\boldsymbol{\sigma}^{\tau}=(\tau\sigma_{x},\sigma_{y})$ the Pauli matrices acting in the A,B sublattice space. In Eq. (1), ${\mathbf{K}}^{\tau}_{1}$ and ${\mathbf{K}}^{\tau}_{2}$ are the Dirac points of valley $\tau$ in the bottom and top layers twisted by angles $\mp\frac{\theta}{2}$. They are equivalent to $\tau{\mathbf{K}}_{1}$ and $\tau{\mathbf{K}}_{2}$ in the moiré Brillouin zone (mBZ), as shown in Fig. 1(c). The interlayer tunneling between the Dirac states in the two layers is described by the matrix
$\begin{split}U=&\left(\begin{array}[]{cc}u_{0}&u_{1}\\ u_{1}&u_{0}\end{array}\right)+\left(\begin{array}[]{cc}u_{0}&u_{1}e^{-i\tau\frac{2\pi}{3}}\\ u_{1}e^{i\tau\frac{2\pi}{3}}&u_{0}\end{array}\right)e^{-i\tau{\mathbf{G}}_{1}\cdot{\mathbf{r}}}\\ &+\left(\begin{array}[]{cc}u_{0}&u_{1}e^{i\tau\frac{2\pi}{3}}\\ u_{1}e^{-i\tau\frac{2\pi}{3}}&u_{0}\end{array}\right)e^{-i\tau({\mathbf{G}}_{1}+{\mathbf{G}}_{2})\cdot{\mathbf{r}}}\end{split}$ (2)
where $u_{0}$ and $u_{1}$ are the intra- and inter-sublattice interlayer tunneling amplitudes. We adopt realistic parameters for the continuum model: $\hbar v_{F}/a_{0}=2.365$ eV, inter-sublattice interlayer coupling $u_{1}=0.11$ eV, and intra-sublattice interlayer coupling $u_{0}=0.06$ eV. The twist angle is fixed at the magic value $\theta$=1.086, at which the low-energy moiré bands become perfectly flat in the limit $u_{0}\to 0$ Tarnopolsky et al. (2019). The low-energy flat bands exhibit a bandwidth of approximately 2 meV and are well separated from the higher-energy remote bands by an energy gap of about 50 meV as shown in Fig. 1. The eigenstates of the continuum Hamiltonian $H_{BM}^{\tau}$ have the Bloch wavefunction form
$\psi_{m,\tau,{\mathbf{k}}}^{X}({\mathbf{r}})=\sum_{{\mathbf{G}}}u_{m,\tau;{\mathbf{G}},X}({\mathbf{k}})e^{i({\mathbf{k}}+{\mathbf{G}})\cdot{\mathbf{r}}}$ (3)
where $X=\{A_{1},B_{1},A_{2},B_{2}\}$ is the layer and sublattice index with the eigen-band energy $\epsilon_{m{\mathbf{k}}\tau}$ and $\mathbf{G}$} is the general reciprocal lattice vector in the mBZ. Here, $m$ and $\tau$ are the band and valley indices, and the spin index $s$ is omitted since the Hamiltonian is spin independent. In the calculation, we take 121 $\mathbf{G}$} vectors as expressed as ${\mathbf{G}}=n_{1}{\mathbf{G}}_{1}+n_{2}{\mathbf{G}}_{2}$ with $n_{1},n_{2}\in[-5,5]$, which is enough to produce the flat bands.
Figure 1: (a) Low-energy band structure of non-interacting TBG calculated using the BM model. The black solid and red dashed curves correspond to the $K^{\prime}$ and $K$ valleys, respectively. (b) Zoomed-in view of the isolated flat bands near charge neutrality, highlighting their separation from the higher-energy remote bands. (c) Schematic illustration of Brillouin zone folding in TBG. The small hexagon denotes the moiré Brillouin zone (mBZ), reciprocal to the moiré superlattice. The cyan arrows indicate the momentum path in the enlarged moiré Brillouin zone along which the band dispersions in panels (a) and (b) are plotted.
III Interacting case
The full Hamiltonian is written as $H=H_{0}+H_{int}$, with the non-interacting part given by
$H_{0}=\sum_{\tau}H_{BM}^{\tau}=\sum_{m{\mathbf{k}}s\tau}\epsilon_{m{\mathbf{k}}\tau}d_{{\mathbf{k}},m,s,\tau}^{\dagger}d_{{\mathbf{k}},m,s,\tau},$ (4)
Here, $d^{\dagger}_{{\mathbf{k}},m,s,\tau}$ and $d_{{\mathbf{k}},m,s,\tau}$ are the creation and annihilation operators in the eigen-band basis, respectively. To incorporate interactions, we project the single-gate-screened Coulomb interaction onto the two lowest-energy flat bands for each spin and valley flavor, and solve the resulting problem self-consistently within a standard Hartree-Fock (HF) approximation in momentum space. Within this treatment, interaction effects are retained only in the active flat-band subspace, while the remote bands are kept fixed. The detailed construction of the projected interaction, the form factors, and the resulting HF Hamiltonian is presented in the Supplemental Material sm , following the standard formulations of Refs. Bernevig et al. (2021); Kang et al. (2021); Zhang et al. (2020). In the numerical calculations, we adopt a dielectric constant $\varepsilon=7$ and a gate distance $d_{s}=40$ nm for the single gate screened Coulomb interaction. In the HF calculation, we take $N_{k}=30\times 30=900$ kpoints in the mBZ.
It is well established that electron–electron Coulomb interactions in twisted bilayer graphene can drive spontaneous symmetry breaking, giving rise to a series of correlated insulating states at commensurate integer fillings. Among these, states that break the combined twofold-rotation and time-reversal symmetry $C_{2}\mathcal{T}$, where $\mathcal{T}$ denotes time reversal, are energetically competitive. In particular, at fillings $\nu=\pm 3$, the TRSB Chern insulating states can emerge as ground states even in the absence of an explicit staggered sublattice potential. Within this projected-basis HF framework, we systematically identify the correlated insulating states at all integer fillings and compute the orbital magnetization of these interaction-driven phases.
IV Orbital magnetization
The topology of the correlated electronic states and their response to an external magnetic field are governed by the Berry curvature and orbital magnetization of the interaction-renormalized bands. In our approach, the Coulomb interaction is projected onto the lowest two flat bands per spin and per valley. As a consequence, only those HF renormalized flat bands that break the combined $C_{2}\mathcal{T}$ symmetry can acquire nontrivial Berry curvature and hence finite Chern numbers. Since the Berry curvature depends only on the periodic part of the Bloch wave functions, it can be directly computed from the HF eigenstates as
${\bf\Omega}_{n}({\mathbf{k}})=i\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}^{HF}(\bm{k})\middle|\partial_{\nu}u_{n}^{HF}(\bm{k})\right\rangle,$ (5)
where $n$ denotes the band index, $\epsilon_{\mu\nu}$ is the antisymmetric tensor, $\mu,\nu$ are Cartesian indices, and $\partial_{\mu}\equiv\partial/\partial k_{\mu}$. For the orbital magnetization, it has been recently proved Kang et al. (2025) that both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ can be evaluated using the HF Hamiltonian and its eigenfunctions Souza and Vanderbilt (2008); Resta (2020); Ji et al. (2025):
$M_{orb}=\frac{e}{2i\hbar}\sum_{n\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\left|\left(H^{\rm HF}(\bm{k})+E_{n}^{\rm HF}(\bm{k})-2\mu\right)\right|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(E_{n}^{\rm HF}(\bm{k})-\mu\right)$ (6)
$m_{SR}=\frac{e}{2i\hbar}\sum_{n,\alpha}\sum_{\bm{k}}\sum_{\mu\nu}\epsilon_{\mu\nu}\left\langle\partial_{\mu}u_{n}(\bm{k})\middle|u_{\alpha}(\bm{k})\right\rangle\left(E^{\rm HF}_{\alpha}(\bm{k})-E_{n}^{\rm HF}(\bm{k})\right)\left\langle u_{\alpha}(\bm{k})\middle|\partial_{\nu}u_{n}(\bm{k})\right\rangle n_{F}\left(E_{n}^{\rm HF}(\bm{k})-\mu\right)\left(1-n_{F}\left(E_{\alpha}^{\rm HF}(\bm{k})-\mu\right)\right)$ (7)
where $n_{F}(E)=\Theta(\mu-E)$ is the Fermi–Dirac distribution function at $T=0$ and $\mu$ is the chemical potential. These quantities can be reformulated in terms of projection operators.
For orbital magnetization, it is important to emphasize that the Hartree-Fock Hamiltonian $H^{\mathrm{HF}}(\mathbf{k})$ entering Eqs. 6 and 7 is, strictly speaking, defined in the full Hilbert space, including the active flat-band sector together with both the occupied and empty remote bands. The derivation in Ref. Kang et al. (2025) assumes that the HF problem is solved self-consistently in this full Hilbert space. In the present work, however, the HF calculation is carried out only within the active flat-band subspace, while the remote bands are kept fixed. This frozen-band treatment may therefore introduce quantitative errors into the resulting orbital magnetization. A fully self-consistent $B=0$ HF calculation including additional remote bands, and a careful study of the convergence of the orbital magnetization with respect to that extension, are left for future work.
Even within this approximation, Eqs. 6 and 7 still involve sums over infinitely many occupied and empty remote states. To make the calculation tractable, we evaluate $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ in a truncated Hilbert space. For clarity of notation, we label the HF Bloch states $|u_{n}^{A/R+/R-}(\mathbf{k})\rangle$ by the superscripts $A$, $R+$, and $R$-, which denote the active bands, the empty remote bands, and the occupied remote bands, respectively. The remote bands are indexed outward from charge neutrality, such that the states $|u_{n}^{R+}(\mathbf{k})\rangle$ and $|u_{n}^{R-}(-\mathbf{k})\rangle$ are related by particle-hole symmetry Bernevig et al. (2021). On this basis, we introduce two projection operators onto the truncated occupied and empty sectors:
$P_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{occupied}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R-}_{n}(\mathbf{k})\rangle\langle u^{R-}_{n}(\mathbf{k})|$ (8)
$Q_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{empty}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R+}_{n}(\mathbf{k})\rangle\langle u^{R+}_{n}(\mathbf{k})|,$ (9)
These two projection operators for truncated occupied and empty states explicitly depend on the choice of $n_{\rm cut}$, and thus $P_{n_{\rm cut}}(k)+Q_{n_{\rm cut}}(k)=I$ only when $n_{\rm cut}\rightarrow\infty$. In this representation, we introduce the orbital magnetization for the truncated Hilbert space that is expressed as sm
$M^{\rm trun}_{orb}=-\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W^{\rm trun}_{xy}(\bm{k})-N^{\rm trun}_{xy}(\bm{k})\right]$ (10)
$m^{\rm trun}_{SR}=\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W^{\rm trun}_{xy}(\bm{k})+N^{\rm trun}_{xy}(\bm{k})\right],$ (11)
where
$W^{\rm trun}_{\mu\nu}(\bm{k})=\mathrm{Tr}\left[P_{n_{cut}}(\bm{k})\big(\partial_{\mu}P_{n_{cut}}(\bm{k})\big)\big(\partial_{\nu}Q_{n_{cut}}(\bm{k})\big)\big(H^{HF}(\bm{k})-\mu\big)\right]$ (12)
and
$N^{\rm trun}_{\mu\nu}(\bm{k})=\mathrm{Tr}\left[Q_{n_{cut}}(\bm{k})\big(\partial_{\mu}P_{n_{cut}}(\bm{k})\big)\big(\partial_{\nu}Q_{n_{cut}}(\bm{k})\big)\big(H^{HF}(\bm{k})-\mu\big)\right].$ (13)
Because the projection operators $P_{n_{cut}}(\mathbf{k})$ and $Q_{n_{cut}}(\mathbf{k})$ are manifestly gauge invariant, this formulation avoids the gauge ambiguities that typically arise in direct evaluations of orbital magnetization based on $\bm{k}$-derivatives of Bloch wave functions. Moreover, it circumvents the singularities associated with band degeneracy points. Formally, physical quantities are defined in the limit $n_{\mathrm{cut}}\to\infty$. Nevertheless, we anticipate rapid convergence with increasing $n_{\mathrm{cut}}$, enabling reliable numerical evaluation within a finite truncated Hilbert space.
IV.1 Valley magnetization for the BM model
We first apply the projection-matrix formalism to the non-interacting BM model. To generate a finite valley magnetization, we introduce a staggered potential through an additional term $\Delta_{C2B}\sigma_{z}$ in the BM Hamiltonian. This term breaks $C_{2}$ symmetry and thus gaps the Dirac points, as shown in Fig. 2(a). Because the staggered potential preserves time-reversal symmetry, the orbital magnetizations in the two valleys remain equal in magnitude and opposite in sign, such that the total orbital magnetization vanishes. We therefore restrict our analysis to the $\tau=$- valley and evaluate $M_{\mathrm{orb}}$ within this sector. The self-rotation contribution $m_{\mathrm{SR}}$ is not evaluated in this case, as it is logarithmic divergent in the ultraviolet limit unless an additional high-energy cutoff is introduced sm .
To highlight the importance of the truncation procedure, we compare our scheme with an alternative prescription, denoted m1, in which only the occupied projection matrix $P_{n_{\rm cut}}(\mathbf{k})$ is truncated, while the empty-state projection matrix is taken as $Q_{n_{\rm cut}}^{\prime}(\mathbf{k})=I-P_{n_{\rm cut}}(\mathbf{k})$. As shown in Fig. 2(b), truncating both projection matrices yields smooth and rapid convergence of $M_{\mathrm{orb}}$ with increasing $n_{\mathrm{cut}}$ (black curves). In contrast, scheme m1 exhibits substantial oscillations and attains convergence only when essentially all remote bands are included in the occupied projection matrix, where the two prescriptions become equivalent. The origin of this oscillatory behavior and the numerical stability of the symmetric truncation scheme are discussed in the Supplemental Material sm .
Figure 2: (a) Band structure for valley ($\tau=$-) in the non-interacting BM model with a staggered potential $\Delta_{C2B}=20$ meV. The gapped low-energy flat bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). (b) Calculated total orbital magnetization $M_{\mathrm{orb}}$ per moiré cell for the valley - state shown in (a), obtained using two truncation schemes within the projection-matrix formalism and plotted as functions of the number of remote-band pairs $n_{\mathrm{cut}}$. When evaluating $M_{\mathrm{orb}}$, the chemical potential is fixed at $\mu=-10$ meV, inside the gap. The inset shows the enlarged view of panel (b), highlighting the oscillatory behavior of scheme m1.
IV.2 $\nu=\pm 3$
With the truncation scheme validated in the non-interacting BM model, we now apply the same formalism to the interacting Hartree-Fock states, beginning with the Chern insulating phases at $\nu=\pm 3$. The self-consistent Hartree–Fock calculation yields a spin–valley polarized Chern insulating ground state characterized by spontaneous breaking of the combined $C_{2}\mathcal{T}$ symmetry. The corresponding band dispersions, shown in Fig. 3(a) and (b), indicate that among the eight active bands, a single pair associated with one spin–valley flavor develops a gap due to $C_{2}\mathcal{T}$ symmetry breaking and acquires a finite Chern number. The remaining three pairs of bands, corresponding to the other spin–valley flavors (shown in black), remain topological trivial. The symmetry breaking generates an indirect energy gap of approximately 13 meV between the occupied Chern band and the higher-energy bands. These states correspond to the experimentally observed quantum anomalous Hall phases with Chern number $C=1$ (C=1C=-1) at fillings $\nu=3$ ($\nu=-3$Sharpe et al. (2019); Serlin et al. (2020); Nuckolls et al. (2020).
Figure 3: (a,b) Band dispersions of the Chern insulating states at fillings $\nu=3$ (a) and $\nu=-3$ (b), obtained within the HF approximation. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands that are excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1), corresponding to the experimentally observed Chern insulating states with $C=\pm 1$ at $\nu=\pm 3$. (c,d) Calculated total orbital magnetization $M_{\mathrm{orb}}$ and self-rotation contribution $m_{\mathrm{SR}}$ per moiré cell for the states shown in (a,b), plotted as functions of the number of remote-band pairs ($n_{\mathrm{cut}}$) included in the projection-matrix formalism. Both quantities exhibit clear convergence for $n_{\mathrm{cut}}\approx 20$. When evaluating $M_{\mathrm{orb}}$, the chemical potential $\mu$ is fixed at the top of the valence band for $\nu=3$ and at the bottom of the conduction band for $\nu=-3$, as indicated by the horizontal dashed lines in panels (a) and (b).
We next evaluate the orbital magnetization of these states using Eqs. 10 and 11. As shown in Fig. 3(c,d), both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ converge by $n_{\mathrm{cut}}\approx 20$. The results also respect the particle-hole symmetry relating fillings $\nu=3$ and $\nu=-3$, with $M_{\mathrm{orb}}(\nu=3)=M_{\mathrm{orb}}(\nu=-3)$ and $m_{\mathrm{SR}}(\nu=3)=-m_{\mathrm{SR}}(\nu=-3)$ sm , consistent with the experimentally observed Landau-fan structure Das et al. (2021); Nuckolls et al. (2020). Since convergence is already achieved for $n_{\mathrm{cut}}\approx 20$, we use $n_{\mathrm{cut}}=30$ in the following calculations to ensure numerical stability.
One might be tempted to conclude that both the orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ arise exclusively from the occupied Chern bands that explicitly break time-reversal symmetry, and that only these bands need to be considered, in analogy with the case of Berry curvature. Such reasoning, however, is not generally valid. Although the occupied states of the remote bands, $|u_{n}(\mathbf{k})\rangle$, remain invariant under time-reversal symmetry, their momentum derivatives $|\partial_{\mu}u_{n}(\mathbf{k})\rangle$ generally possess non-vanishing overlaps with those states in the active bands. Consequently, once time-reversal symmetry is broken within the active bands, the nominally time-reversal-symmetric remote bands acquire finite contributions to both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$. The need to include at least $20$ remote-band pairs to reach numerical convergence provides direct evidence that these contributions are substantial.
To make the role of the remote bands more transparent, we fix the HF band structure obtained at $\nu=3$ and introduce an auxiliary energy level $\mu^{\prime}$ in place of the physical chemical potential $\mu$. We then vary $\mu^{\prime}$ rigidly from 40-40 meV, well below the active HF bands, to $80$ meV, well above them, and compute both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ as functions of $\mu^{\prime}$. For a fixed HF spectrum, the resulting $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ represent the contributions from all active bands below $\mu^{\prime}$ together with the negative-energy remote bands. As shown in Fig. 4(a,b), the blue plateaus in both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ for $\mu^{\prime}\lesssim-20$ meV arise entirely from the remote bands and reach magnitudes as large as $\sim 7.5\mu_{B}$, directly demonstrating their sizable contribution to the orbital magnetization. In this regime, $M_{\mathrm{orb}}=m_{\mathrm{SR}}$ because the remote bands are topologically trivial and carry zero Berry curvature.
More physically relevant is the behavior of $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ within the Chern insulating gap, indicated by the red dots in Fig. 4(a,b). Inside the Chern gap, the auxiliary energy level $\mu^{\prime}$ is equivalent to the true chemical potential $\mu$ of the Chern insulating state. The total orbital magnetization $M_{\mathrm{orb}}$ exhibits the expected linear dependence on $\mu$. Thermodynamics gives the relation $\partial M_{orb}/\partial\mu=\partial N/\partial B$, consistent with the Diophantine equation that characterizes quantum Hall states. Taking the derivative of Eq.6 with respect to $\mu$ yields
$\frac{\partial M_{orb}}{\partial\mu}=S\frac{e}{h}C,$ (14)
where $C$ is the Chern number. The linear dependence leads to a sign reversal of $M_{\mathrm{orb}}$ across the gap, corresponding to magnetization reversal previously analyzed in the non-interacting case Zhu et al. (2020). The in-gap plateau of $m_{\mathrm{SR}}$ demonstrates that the Chern insulating state possesses a large self-rotation contribution of approximately $40\mu_{B}$, which is directly accessible through magneto-optical probes such as magnetic circular dichroism.
Figure 4: Calculated total orbital magnetization $M_{\mathrm{orb}}$ (a) and self-rotation contribution $m_{\mathrm{SR}}$ (b) per moiré cell as functions of the energy level $\mu^{\prime}$ for the Chern insulating state at $\nu=3$ inside the insulating gap, whose band dispersion is shown in Fig. 3(a). The red dots denote the values within the Chern insulating gap, where $M_{\mathrm{orb}}$ shows a linear dependence with $\mu^{\prime}$($\mu$) and changes sign within the gap while $m_{\mathrm{SR}}$ remains constant. The blue dots represent the contribution from the remote bands arising from the time-reversal symmetry breaking of the HF Hamiltonian $H_{\mathrm{tot}}^{\mathrm{HF}}$. The number of remote band pairs $n_{cut}$ is set to 30 in the calculation.
IV.3 Competing states at $\nu=\pm 1$
We next consider the correlated insulating states at fillings $\nu=\pm 1$. Owing to the particle–hole symmetry relating $\nu$ and $\nu$, we focus on the case $\nu=1$. The Hartree–Fock calculation yields two types of nearly degenerate correlated insulating states. The lowest-energy state is a mixed phase combining intervalley coherence (IVC) and $C_{2}\mathcal{T}$ symmetry breaking. Its band dispersion is shown in Fig. 5(b). In this state, the five valence bands consist of two bands associated with a spin-polarized IVC sector carrying zero Chern number, one $C_{2}\mathcal{T}$-broken band with finite Chern number, and an additional pair of ungapped bands from a single spin-valley flavor, yielding a total Chern number $C=1$ Zhang et al. (2020). This state is lower in energy by approximately 1.52 meV per moiré cell than a family of states in which three valence bands break $C_{2}\mathcal{T}$ symmetry and each carry a finite Chern number. We refer to the latter as correlated Chern insulators (CCIs). Among these degenerate CCI solutions, we present the state with the largest Chern number, namely $C=3$, whose dispersion is shown in Fig. 5(a).
We then evaluate the orbital magnetization for both states. As shown in Fig. 5(c), the total orbital magnetization $M_{\mathrm{orb}}$ depends linearly on the chemical potential within the insulating gap and changes sign across the gap for both states. Both the magnitude and the slope of $M_{\mathrm{orb}}$ are larger for the CCI than for the mixed IVC state, reflecting its larger Chern number. This implies that the CCI is energetically favored in the presence of a finite magnetic field, consistent with experimental observations. The self-rotation contribution $m_{\mathrm{SR}}$, shown in Fig. 5(d), is likewise significantly larger in magnitude for the CCI than for the mixed IVC state. Using the difference in orbital magnetization at the center of the gap, $\delta M_{\mathrm{orb}}=4.31\,\mu_{B}$ per moiré cell, together with the energy difference between the two insulating states, $\delta E=1.52$ meV per moiré cell, we estimate a critical magnetic field for the transition from the mixed IVC state to the CCI of $B_{c}\sim\delta E/\delta M_{\mathrm{orb}}\approx 6.1$ T. We also note a sign-convention subtlety in comparison with experiment. In Landau-fan analyses based on the Streda formula, the Chern number is extracted from the slope of the carrier density with respect to the applied magnetic field. Because the charge carriers are electrons and therefore carry negative charge, the physical Chern number is the negative of this slope, in agreement with the convention used in our theoretical analysis. Thus, for direct comparison with the experimentally reported values, one should consider the time-reversed partner of the ground state obtained in our calculation, under which both the Chern number and the associated orbital magnetization change sign.
Figure 5: (a,b) HF band dispersions of the correlated Chern insulating state (CCI) (a) and the mixed intervalley-coherent (IVC) state (b) at filling $\nu=1$. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). For the CCI, the Chern bands are triply degenerate, yielding a total Chern number $C=3$. (c,d) Total orbital magnetization $M_{\mathrm{orb}}$ (c) and self-rotation contribution $m_{\mathrm{SR}}$ (d) per moiré cell as functions of the chemical potential $\mu$ inside the insulating gap for the two states shown in (a) and (b). Solid circles correspond to the CCI, and solid squares to the mixed IVC state. $M_{\mathrm{orb}}$ varies linearly with $\mu$ and changes sign across the gap, while $m_{\mathrm{SR}}$ remains constant. The number of remote-band pairs is fixed at $n_{\mathrm{cut}}=30$ in the calculation.
V Conclusion
In this work, we developed a practical and gauge-invariant framework to compute orbital magnetization in correlated states of magic-angle twisted bilayer graphene within self-consistent Hartree–Fock theory. Based on the projector formulation of the recently developed theory of orbital magnetization in the interacting system Kang et al. (2025), we evaluated both the total orbital magnetization $M_{\mathrm{orb}}$ and the self-rotation contribution $m_{\mathrm{SR}}$ directly from the HF Hamiltonian and its eigenstates, thereby avoiding gauge ambiguities associated with Bloch-wavefunction derivatives. A key advantage of this approach is that it enables controlled inclusion of remote-band effects through the projection matrices. We demonstrated that, unlike the Chern number, the orbital magnetization is highly sensitive to remote bands and requires careful convergence with respect to the number of remote band pairs $n_{\mathrm{cut}}$, which saturates for $n_{\mathrm{cut}}\gtrsim 20$ in our calculations.
Applying this method to integer fillings, we obtained converged orbital magnetization for the TRSB Chern insulating states at $\nu=\pm 3$ and for competing correlated phases at $\nu=1$. Our results ensure the thermodynamic linear dependence of $M_{\mathrm{orb}}$ on the chemical potential inside the insulating gaps with the slope determined by the Chern number, while yielding a sizable and constant in-gap $m_{\mathrm{SR}}$ relevant for magneto-optical probes. More broadly, the gauge-invariant projector formalism introduced here provides a systematic route to quantify orbital magnetization in interaction-driven moiré phases and can be readily extended to other moiré materials.
ACKNOWLEDGMENTS
We thank Hui-Ke Jin, Jie Wang, Minxuan Wang, Zhida Song, Wen-Yu He, and especially Oskar Vafek for helpful discussions. J. K. acknowledges the support from the NSFC (Grant No. 12574159 and No. 12074276), the Double First-Class Initiative Fund of ShanghaiTech University, and the start-up grant of ShanghaiTech University. YZ is supported in part by the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800) and National Natural Science Foundation of China (NSFC) Grants No. 12274279. ZW is supported by the U.S. Department of Energy, Basic Energy Sciences, Grant No. DE FG02-99ER45747.
Supplemental Material: Contribution of remote bands to orbital magnetization in twisted bilayer graphene
I HF approximation in the projected basis
In this section, we summarize the standard projected-basis Hartree–Fock formulation used in the main text. We project the Coulomb interaction onto the two lowest-energy flat bands per spin and per valley and work directly in momentum space, following Refs. Bernevig et al. (2021); Kang et al. (2021); Zhang et al. (2020). The interacting part of the Hamiltonian is written as
$H_{int}=\frac{1}{2S}\sum_{{\mathbf{q}}\neq 0}V({\mathbf{q}})\delta\rho_{{\mathbf{q}}}\delta\rho_{-{\mathbf{q}}}\ ,$ (S1)
where $S$ is the total area and $\delta\rho_{{\mathbf{q}}}=\rho_{{\mathbf{q}}}-\bar{\rho}_{{\mathbf{q}}}$ is the density operator relative to charge neutrality of the decoupled bilayer graphene in momentum space. Compared with the normal-ordered interaction supplemented by double-counting subtraction schemes used in some previous HF studies Bultinck et al. (2020); Liu and Dai (2021); Zhang et al. (2020), the interaction in Eq. S1 partially incorporates the renormalization and screening effects of the remote bands on the active flat bands Kang et al. (2021); Bernevig et al. (2021). The density operator can be expanded in the eigen-basis of $H_{0}$ as
$\rho_{{\mathbf{q}}}=\sum_{s,\tau}\sum_{{\mathbf{k}}\in\text{mBZ},m_{1},m_{2}}\lambda_{m_{1},m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})d^{\dagger}_{{\mathbf{k}},m_{1},s,\tau}d_{{\mathbf{k}}+{\mathbf{q}},m_{2},s,\tau}\,$ (S2)
with the quantum-averaged background density
$\bar{\rho}_{{\mathbf{q}}}=\sum_{{\mathbf{G}},\tau,m}\delta_{{\mathbf{q}},{\mathbf{G}}}\sum_{{\mathbf{k}}\in\text{mBZ}}\lambda_{m,m;\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{G}})\ ,$ (S3)
where the form factor is defined by
$\begin{split}\lambda_{m_{1}m_{2};\tau}({\mathbf{k}},{\mathbf{k}}+{\mathbf{q}})&=\left<u_{m_{1},\tau}({\mathbf{k}})|u_{m_{2},\tau}({\mathbf{k}}+{\mathbf{q}})\right>=\sum_{{\mathbf{G}}^{{}^{\prime}},X}u^{*}_{m_{1},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}})u_{m_{2},\tau;{\mathbf{G}}^{{}^{\prime}},X}({\mathbf{k}}+{\mathbf{q}}).\end{split}$ (S4)
Here, $\mathbf{k}$} is restricted to the mBZ while $\mathbf{q}$} can extend to the original Brillouin zone of monolayer graphene. The operator $d^{\dagger}$ obeys the periodic condition $d^{\dagger}_{{\mathbf{k}},m,s,\tau}=d^{\dagger}_{{\mathbf{k}}+{\mathbf{G}},m,s,\tau}$, which is equivalent to the Bloch-wavefunction relation $u_{m,\tau;{\mathbf{G}},X}({\mathbf{k}}+{\mathbf{G}}_{0})=u_{m,\tau;{\mathbf{G}}+{\mathbf{G}}_{0},X}({\mathbf{k}})$. We take $V({\mathbf{q}})$ to be the single-gate-screened Coulomb potential Bultinck et al. (2020),
$V({\mathbf{q}})=\frac{e^{2}}{2\varepsilon\varepsilon_{0}q}(1-e^{-2qd_{s}})\ ,$ (S5)
with $\varepsilon=7$ and $d_{s}=40$ nm.
Within this projected basis, the quartic interaction is decoupled in the standard Hartree and Fock channels in terms of the single-particle density matrix
$\rho({\mathbf{k}})_{m_{1},s,\tau;m_{2},s^{\prime}\tau^{\prime}}=\langle d^{\dagger}_{{\mathbf{k}},m_{1},s,\tau}d_{{\mathbf{k}},m_{2},s^{\prime},\tau^{\prime}}\rangle.$ (S6)
We then solve the resulting HF Hamiltonian self-consistently at each integer filling. The converged HF eigenvalues and eigenvectors are used in Eqs. 5-11 of the main text to evaluate the Berry curvature, orbital magnetization, and self-rotation contribution.
II Orbital Magnetization for a Gapped Dirac Cone
In this section, we consider a noninteracting two-band toy model describing a gapped Dirac cone, with the Hamiltonian
$H(\bm{k})=\hbar v_{F}\bm{\sigma}\cdot\bm{k}+\Delta\sigma_{z}\ .$ (S7)
The spectrum consists of two states at each $\bm{k}$,
$E_{\pm}=\pm\sqrt{\Delta^{2}+(\hbar v_{F}k)^{2}}\ ,$ (S8)
with corresponding eigenstates
$|u_{+}\rangle$ $=\left(\cos\frac{\theta}{2},\ \sin\frac{\theta}{2}e^{i\phi}\right)^{T},$ (S9)
$|u_{-}\rangle$ $=\left(-\sin\frac{\theta}{2},\ \cos\frac{\theta}{2}e^{i\phi}\right)^{T},$ (S10)
where
$\theta=\tan^{-1}\frac{\hbar v_{F}k}{\Delta}\ ,\qquad\phi=\tan^{-1}\frac{k_{y}}{k_{x}}\ .$ (S11)
Using the identity
$\langle\partial_{\mu}u_{-}|u_{+}\rangle=\frac{\langle u_{-}|\partial_{\mu}H|u_{+}\rangle}{E_{-}-E_{+}}=\hbar v_{F}\frac{\langle u_{-}|\sigma_{\mu}|u_{+}\rangle}{E_{-}-E_{+}}\ ,$ (S12)
we obtain
$\langle\partial_{x}u_{-}|u_{+}\rangle$ $=\hbar v_{F}\frac{\langle u_{-}|\sigma_{x}|u_{+}\rangle}{E_{-}-E_{+}}=-\frac{\hbar v_{F}}{2\sqrt{\Delta^{2}+(\hbar v_{F}k)^{2}}}\left(\cos^{2}\frac{\theta}{2}e^{-i\phi}-\sin^{2}\frac{\theta}{2}e^{i\phi}\right)\ ,$ (S13)
$\langle\partial_{y}u_{-}|u_{+}\rangle$ $=\hbar v_{F}\frac{\langle u_{-}|\sigma_{y}|u_{+}\rangle}{E_{-}-E_{+}}=i\frac{\hbar v_{F}}{2\sqrt{\Delta^{2}+(\hbar v_{F}k)^{2}}}\left(\cos^{2}\frac{\theta}{2}e^{-i\phi}+\sin^{2}\frac{\theta}{2}e^{i\phi}\right)\ .$ (S14)
Substituting these matrix elements into the general expression for the orbital magnetization gives
$M_{\text{orb}}$ $=\frac{e}{2i\hbar}\sum_{\bm{k}}\Theta(\mu-E_{-}(\bm{k}))\Theta(E_{+}(\bm{k})-\mu)\left(\langle\partial_{x}u_{-}|u_{+}\rangle\langle u_{+}|\partial_{y}u_{-}\rangle(E_{+}+E_{-}-2\mu)-\mathrm{c.c.}\right)$
$=-\frac{e\mu}{\hbar}\sum_{\bm{k}}\Theta(\mu-E_{-}(\bm{k}))\Theta(E_{+}(\bm{k})-\mu)\frac{(\hbar v_{F})^{2}}{\Delta^{2}+(\hbar v_{F}k)^{2}}\left(\cos^{4}\frac{\theta}{2}-\sin^{4}\frac{\theta}{2}\right)\ .$ (S15)
$\Longrightarrow\quad\frac{M_{\text{orb}}}{S}$ $=-\frac{e\mu}{\hbar}\int_{0}^{\infty}\frac{k{\rm d}k}{2\pi}\frac{(\hbar v_{F})^{2}}{\Delta^{2}+(\hbar v_{F}k)^{2}}\cos\theta$
$=-\frac{e\mu}{2h}\int_{0}^{\infty}k{\rm d}k\ \Theta\left(\mu+\sqrt{\Delta^{2}+(\hbar v_{F}k)^{2}}\right)\Theta\left(\sqrt{\Delta^{2}+(\hbar v_{F}k)^{2}}-\mu\right)\frac{(\hbar v_{F})^{2}\Delta}{(\Delta^{2}+(\hbar v_{F}k)^{2})^{3/2}}$
$=\left\{\begin{array}[]{ll}-\dfrac{e\mu}{2h}\ ,&|\mu|\leq\Delta\ ,\\[6.0pt] -\dfrac{e\Delta}{2h}\mathrm{sgn}(\mu)\ ,&|\mu|>\Delta\ .\end{array}\right$. (S18)
It is also instructive to evaluate the self-rotation contribution $m_{\text{SR}}$. This yields
$\frac{m_{\text{SR}}}{S}$ $=-\frac{e}{h}\int_{0}^{\infty}k{\rm d}k\ \frac{(\hbar v_{F})^{2}\Delta}{\Delta^{2}+(\hbar v_{F}k)^{2}}\ .$ (S19)
This integral diverges logarithmically in the ultraviolet limit $k\rightarrow\infty$. Within the present continuum description, this indicates that the valley self-rotation contribution $m_{\text{SR}}$ is ultraviolet divergent, and therefore would also diverge in hBN-aligned Dirac systems unless a suitable high-energy cutoff is introduced.
III Projection matrix formalism for orbital magnetization
In this section, we derive the expression for $M_{\text{orb}}$ and $m_{\text{SR}}$ in terms of projection operators. We begin from Eq. 6 and 7, which give the general expression for the orbital magnetization. However, these formulas are not convenient for numerical implementation, since the derivative of the Bloch states is gauge dependent, and the summation over the occupied bands involves a large number of bands for moiré systems.
We first note that the expression of $M_{orb}$ in Eq. 6 can be rewritten as
$M_{\text{orb}}$ $=\frac{e}{2i\hbar}\sum_{n,\bm{k}}n_{F}\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)\epsilon_{\mu\nu}\sum_{\alpha}\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\big(E_{\alpha}^{\text{HF}}(\bm{k})+E_{n}^{\text{HF}}(\bm{k})-2\mu\big)\langle u_{\alpha}(\bm{k})|\partial_{\nu}u_{n}(\bm{k})\rangle$ (S20)
The summation over $\alpha$ can be decomposed into contributions from occupied and empty states. Accordingly, $M_{\text{orb}}$ can be expressed as
$M_{\text{orb}}$ $=M_{\text{orb}}^{(1)}+M_{\text{orb}}^{(2)}$ (S21)
$M_{\text{orb}}^{(1)}$ $=\frac{e}{2i\hbar}\sum_{n,\alpha,\bm{k}}n_{F}\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)n_{F}\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)\epsilon_{\mu\nu}\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\big(E_{\alpha}^{\text{HF}}(\bm{k})+E_{n}^{\text{HF}}(\bm{k})-2\mu\big)\langle u_{\alpha}(\bm{k})|\partial_{\nu}u_{n}(\bm{k})\rangle$ (S22)
$M_{\text{orb}}^{(2)}$ $=\frac{e}{2i\hbar}\sum_{n,\alpha,\bm{k}}n_{F}\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)\left(1-n_{F}\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)\right)\epsilon_{\mu\nu}\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\big(E_{\alpha}^{\text{HF}}(\bm{k})+E_{n}^{\text{HF}}(\bm{k})-2\mu\big)\langle u_{\alpha}(\bm{k})|\partial_{\nu}u_{n}(\bm{k})\rangle$ (S23)
Using the orthogonality relation,
$\langle u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle=\delta_{\alpha n}\ ,$
we obtain
$\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle=-\langle u_{n}(\bm{k})|\partial_{\mu}u_{\alpha}(\bm{k})\rangle\ .$
It then follows that $M_{\text{orb}}^{(1)}$ can be rewritten as
$M_{\text{orb}}^{(1)}$ $=\frac{e}{2i\hbar}\sum_{n,\alpha,\bm{k}}n_{F}\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)n_{F}\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)\epsilon_{\mu\nu}\langle u_{n}(\bm{k})|\partial_{\mu}u_{\alpha}(\bm{k})\rangle\big(E_{\alpha}^{\text{HF}}(\bm{k})+E_{n}^{\text{HF}}(\bm{k})-2\mu\big)\langle\partial_{\nu}u_{\alpha}(\bm{k})|u_{n}(\bm{k})\rangle$
$=\frac{e}{2i\hbar}\sum_{n,\alpha,\bm{k}}n_{F}\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)n_{F}\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)\epsilon_{\mu\nu}\langle\partial_{\nu}u_{\alpha}(\bm{k})|u_{n}(\bm{k})\rangle\big(E_{\alpha}^{\text{HF}}(\bm{k})+E_{n}^{\text{HF}}(\bm{k})-2\mu\big)\langle u_{n}(\bm{k})|\partial_{\mu}u_{\alpha}(\bm{k})\rangle$
$=-M_{\text{orb}}^{(1)}\ .$ (S24)
Therefore,
$M_{\text{orb}}^{(1)}=0\ ,$ (S25)
so that this term does not contribute to the orbital magnetization.
We next rewrite $M_{\mathrm{orb}}^{(2)}$ in terms of projectors. For the full HF Hamiltonian $H^{\mathrm{HF}}(\bm{k})$, we define the projectors onto the occupied and empty subspaces as
$P(\bm{k})$ $=\sum_{n\in\mathrm{occ}}|u_{n}(\bm{k})\rangle\langle u_{n}(\bm{k})|\ ,$ (S26)
$Q(\bm{k})$ $=\sum_{\alpha\in\mathrm{emp}}|u_{\alpha}(\bm{k})\rangle\langle u_{\alpha}(\bm{k})|\ .$ (S27)
We then introduce the following gauge invariant quantities,
$W_{\mu\nu}(\bm{k})$ $=\mathrm{Tr}\left[P(\bm{k})\big(\partial_{\mu}P(\bm{k})\big)\big(\partial_{\nu}Q(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]\ ,$ (S28)
$N_{\mu\nu}(\bm{k})$ $=\mathrm{Tr}\left[Q(\bm{k})\big(\partial_{\mu}P(\bm{k})\big)\big(\partial_{\nu}Q(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]\ .$ (S29)
Since $PQ=0$, one has
$\big(\partial_{\mu}P\big)Q=-P\big(\partial_{\mu}Q\big)\ .$
Moreover, both $P$ and $Q$ commute with $H^{\text{HF}}$. Consequently,
$W_{\mu\nu}(\bm{k})$ $=\mathrm{Tr}\left[P(\bm{k})\big(\partial_{\mu}P(\bm{k})\big)\big(\partial_{\nu}Q(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]=\mathrm{Tr}\left[\big(\partial_{\mu}P(\bm{k})\big)\big(\partial_{\nu}Q(\bm{k})\big)P(\bm{k})\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]$
$=-\mathrm{Tr}\left[\big(\partial_{\mu}P(\bm{k})\big)Q(\bm{k})\big(\partial_{\nu}P(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]$
$=-\sum_{n,\alpha}n_{F}(E_{n}(\bm{k})-\mu)\big(1-n_{F}(E_{\alpha}(\bm{k})-\mu)\big)\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\langle u_{\alpha}(\bm{k})|\partial_{\nu}u_{n}(\bm{k})\rangle\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)$ (S30)
which implies
$\big(W_{\mu\nu}(\bm{k})\big)^{*}=-\sum_{n,\alpha}n_{F}(E_{n}(\bm{k})-\mu)\big(1-n_{F}(E_{\alpha}(\bm{k})-\mu)\big)\langle u_{\alpha}(\bm{k})|\partial_{\mu}u_{n}(\bm{k})\rangle\langle\partial_{\nu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\left(E_{n}^{\text{HF}}(\bm{k})-\mu\right)=W_{\nu\mu}(\bm{k})\ .$ (S31)
Similarly,
$N_{\mu\nu}(\bm{k})$ $=\mathrm{Tr}\left[Q(\bm{k})\big(\partial_{\mu}P(\bm{k})\big)\big(\partial_{\nu}Q(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]=-\mathrm{Tr}\left[\big(\partial_{\mu}Q(\bm{k})\big)P(\bm{k})\big(\partial_{\nu}Q(\bm{k})\big)\big(H^{\mathrm{HF}}(\bm{k})-\mu\big)\right]$
$=-\sum_{n,\alpha}n_{F}(E^{\mathrm{HF}}_{n}(\bm{k})-\mu)\big(1-n_{F}(E^{\mathrm{HF}}_{\alpha}(\bm{k})-\mu)\big)\langle\partial_{\mu}u_{\alpha}(\bm{k})|u_{n}(\bm{k})\rangle\langle u_{n}(\bm{k})|\partial_{\nu}u_{\alpha}(\bm{k})\rangle\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)$
$=-\sum_{n,\alpha}n_{F}(E^{\mathrm{HF}}_{n}(\bm{k})-\mu)\big(1-n_{F}(E^{\mathrm{HF}}_{\alpha}(\bm{k})-\mu)\big)\langle\partial_{\nu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\langle u_{\alpha}(\bm{k})|\partial_{\mu}u_{n}(\bm{k})\rangle\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)$ (S32)
and therefore
$\big(N_{\mu\nu}(\bm{k})\big)^{*}=-\sum_{n,\alpha}n_{F}(E^{\mathrm{HF}}_{n}(\bm{k})-\mu)\big(1-n_{F}(E^{\mathrm{HF}}_{\alpha}(\bm{k})-\mu)\big)\langle u_{\alpha}(\bm{k})|\partial_{\nu}u_{n}(\bm{k})\rangle\langle\partial_{\mu}u_{n}(\bm{k})|u_{\alpha}(\bm{k})\rangle\left(E_{\alpha}^{\text{HF}}(\bm{k})-\mu\right)=N_{\nu\mu}(\bm{k})\ .$ (S33)
Substituting these results into Eq. S23, we obtain
$M_{\mathrm{orb}}=M_{\mathrm{orb}}^{(2)}=-\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W_{xy}(\bm{k})-N_{xy}(\bm{k})\right]\ .$ (S34)
Similarly, the self-rotation $m_{\text{SR}}$ is given by
$m_{\mathrm{SR}}=\frac{e}{\hbar}\sum_{\bm{k}}\mathrm{Im}\left[W_{xy}(\bm{k})+N_{xy}(\bm{k})\right]\ .$ (S35)
These expressions are manifestly gauge invariant since they depend only on the projection operators, and are therefore numerically more stable than formulas involving explicit derivatives of Bloch wave functions.
We further emphasize that the derivation above does not explicitly rely on the identity $P+Q=I$. This observation suggests that $P$ and $Q$ may be replaced by their counterparts defined in the truncated occupied and empty space. More specifically, we therefore introduce the truncated projectors defined in the main text,
$P_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{occupied}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R-}_{n}(\mathbf{k})\rangle\langle u^{R-}_{n}(\mathbf{k})|,$ (S36)
$Q_{n_{\rm cut}}(\mathbf{k})$ $=\sum_{n\in\mathrm{empty}}|u^{\rm A}_{n}(\mathbf{k})\rangle\langle u^{\rm A}_{n}(\mathbf{k})|+\sum_{n\leq n_{\rm cut}}|u^{R+}_{n}(\mathbf{k})\rangle\langle u^{R+}_{n}(\mathbf{k})|.$ (S37)
Here, $A$, $R$-, and $R+$ denote the active bands, the occupied remote bands, and the empty remote bands, respectively. Replacing $P(\bm{k})$ and $Q(\bm{k})$ by $P_{n_{\rm cut}}(\bm{k})$ and $Q_{n_{\rm cut}}(\bm{k})$ in the definitions of $W_{\mu\nu}$ and $N_{\mu\nu}$ yields the truncated quantities $W_{\mu\nu}^{\rm trun}$ and $N_{\mu\nu}^{\rm trun}$, and hence Eqs. 10 and 11 of the main text used in the numerical calculations. The full result is recovered in the limit $n_{\mathrm{cut}}\to\infty$, while in practice we find rapid convergence with increasing $n_{\mathrm{cut}}$, as benchmarked in sections below.
IV Numerical stability of the symmetric truncation scheme
The numerical stability of the present construction follows from the way in which the Hilbert space is partitioned. Specifically, we decompose the exact projectors as
$P(\bm{k})=P_{n_{\rm cut}}(\bm{k})+P_{1}(\bm{k})\ ,\qquad Q(\bm{k})=Q_{n_{\rm cut}}(\bm{k})+Q_{1}(\bm{k})\ ,$ (S38)
where $P_{1}(\bm{k})$ and $Q_{1}(\bm{k})$ denote the omitted occupied and empty remote-band sectors, respectively.
As discussed in the previous section, the orbital magnetization is expressed in terms of the two gauge-invariant quantities $W_{\mu\nu}$ and $N_{\mu\nu}$ defined in Eqs. S28 and S29. In moiré systems, however, the large number of occupied and empty bands leads to a substantial computational cost for calculations in the full Hilbert space. One therefore replaces these projectors by their truncated counterparts, and correspondingly replaces $W_{\mu\nu}$ and $N_{\mu\nu}$ by the truncated quantities $W_{\mu\nu}^{\text{trun}}$ and $N_{\mu\nu}^{\text{trun}}$ defined in Eqs. 12 and 13.
Following the formalism introduced in the previous section, one readily obtains
$W^{\rm trun}_{\mu\nu}(\bm{k})$ $=-\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}\alpha\leq n_{\rm cut}\\ \alpha\in\rm emp\end{subarray}}\langle\partial_{\mu}u_{n}|u_{\alpha}\rangle\langle u_{\alpha}|\partial_{\nu}u_{n}\rangle\left(E_{n}^{\rm HF}-\mu\right)=-\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}\alpha\leq n_{\rm cut}\\ \alpha\in\rm emp\end{subarray}}\frac{\langle u_{n}|\partial_{\mu}H^{\rm HF}|u_{\alpha}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\frac{\langle u_{\alpha}|\partial_{\nu}H^{\rm HF}|u_{n}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\left(E^{\rm HF}_{n}-\mu\right)$ (S39)
$N^{\rm trun}_{\mu\nu}(\bm{k})$ $=-\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}\alpha\leq n_{\rm cut}\\ \alpha\in\rm emp\end{subarray}}\langle\partial_{\nu}u_{n}|u_{\alpha}\rangle\langle u_{\alpha}|\partial_{\mu}u_{n}\rangle\left(E^{\rm HF}_{\alpha}-\mu\right)=-\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}\alpha\leq n_{\rm cut}\\ \alpha\in\rm emp\end{subarray}}\frac{\langle u_{n}|\partial_{\nu}H^{\rm HF}|u_{\alpha}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\frac{\langle u_{\alpha}|\partial_{\mu}H^{\rm HF}|u_{n}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\left(E^{\rm HF}_{\alpha}-\mu\right)\ ,$ (S40)
where the explicit $\bm{k}$ dependence of the Hamiltonian, the Bloch states, and the corresponding energies has been omitted for notational convenience. In deriving these expressions, we have also used
$\left\langle u_{\alpha}\middle|\partial_{\mu}u_{n}\right\rangle=\frac{\left\langle u_{\alpha}\middle|\partial_{\mu}H^{\mathrm{HF}}\middle|u_{n}\right\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\ .$ (S41)
In the symmetric truncation scheme adopted in this work, the matrix elements entering $W_{\mu\nu}^{\rm trun}$ and $N_{\mu\nu}^{\rm trun}$ always connect a retained occupied state $|u_{n}\rangle$ belonging to $P_{n_{\rm cut}}$ to a retained empty state $|u_{\alpha}\rangle$ belonging to $Q_{n_{\rm cut}}$. These two states lie on opposite sides of the insulating gap, as illustrated schematically in Fig. S1. The denominator $E_{n}^{HF}-E_{\alpha}^{HF}$ is therefore bounded below by the direct gap, which in turn guarantees smooth convergence of the truncated sums despite the formally infinite remote-band sector.
We now examine the difference between the truncated expressions and those defined in the full Hilbert space. Using the relations $PQ=0$ and the fact that both projectors commute with the Hamiltonian, we obtain
$W_{\mu\nu}-W_{\mu\nu}^{\text{trun}}$ $=-\mathrm{Tr}\left\{\left[\big(\partial_{\mu}P_{n_{\text{cut}}}+\partial_{\mu}P_{1}\big)\big(Q_{n_{\text{cut}}}+Q_{1}\big)\big(\partial_{\nu}P_{n_{\text{cut}}}+\partial_{\nu}P_{1}\big)-\big(\partial_{\mu}P_{n_{\text{cut}}}\big)Q_{n_{\text{cut}}}\big(\partial_{\nu}P_{n_{\text{cut}}}\big)\right]\big(H^{\mathrm{HF}}-\mu\big)\right\}$ (S42)
where the explicit $\bm{k}$ dependence has been omitted for notational convenience. Furthermore, one finds that
$\mathrm{Tr}\left[(\partial_{\mu}P_{n_{\text{cut}}})Q(\partial_{\nu}P_{1}\big)\big(H^{\mathrm{HF}}-\mu\big)\right]=\mathrm{Tr}\left[(\partial_{\mu}P_{1})Q(\partial_{\nu}P_{n_{\text{cut}}}\big)\big(H^{\mathrm{HF}}-\mu\big)\right]=0\ .$ (S43)
It then follows that
$W_{\mu\nu}-W_{\mu\nu}^{\text{trun}}$ $=-\mathrm{Tr}\left\{\left[\big(\partial_{\mu}P_{1}\big)Q\big(\partial_{\nu}P_{1}\big)+\big(\partial_{\mu}P_{n_{\text{cut}}}\big)Q_{1}\big(\partial_{\nu}P_{n_{\text{cut}}}\big)\right]\big(H^{\mathrm{HF}}-\mu\big)\right\}$
$=-\left(\sum_{\begin{subarray}{c}n>n_{{\text{cut}}}\\ n\in\rm occ\end{subarray}}\sum_{\alpha\in\text{emp}}+\sum_{\begin{subarray}{c}n\leq n_{{\text{cut}}}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}\alpha>n_{\text{cut}}\\ \alpha\in\text{emp}\end{subarray}}\right)\frac{\left\langle u_{n}\middle|\partial_{\mu}H^{\mathrm{HF}}\middle|u_{\alpha}\right\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\frac{\left\langle u_{\alpha}\middle|\partial_{\nu}H^{\mathrm{HF}}\middle|u_{n}\right\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}}\left(E^{\rm HF}_{n}-\mu\right)\ .$ (S44)
Similarly, for $N_{\mu\nu}$ we obtain
$N_{\mu\nu}-N_{\mu\nu}^{\text{trun}}$ $=-\mathrm{Tr}\left\{\left[\big(\partial_{\nu}Q_{1}\big)P\big(\partial_{\mu}Q_{1}\big)+\big(\partial_{\nu}Q_{n_{\text{cut}}}\big)P_{1}\big(\partial_{\mu}Q_{n_{\text{cut}}}\big)\right]\big(H^{\mathrm{HF}}-\mu\big)\right\}$
$=-\left(\sum_{\begin{subarray}{c}\alpha>n_{{\text{cut}}}\\ \alpha\in\rm emp\end{subarray}}\sum_{n\in\text{occ}}+\sum_{\begin{subarray}{c}\alpha\leq n_{{\text{cut}}}\\ \alpha\in\rm emp\end{subarray}}\sum_{\begin{subarray}{c}n>n_{\text{cut}}\\ n\in\text{occ}\end{subarray}}\right)\frac{\left\langle u_{\alpha}\middle|\partial_{\nu}H^{\mathrm{HF}}\middle|u_{n}\right\rangle}{E^{\rm HF}_{\alpha}-E^{\rm HF}_{n}}\frac{\left\langle u_{n}\middle|\partial_{\mu}H^{\mathrm{HF}}\middle|u_{\alpha}\right\rangle}{E^{\rm HF}_{\alpha}-E^{\rm HF}_{n}}\left(E^{\rm HF}_{\alpha}-\mu\right)\ .$ (S45)
In the above expressions for differences, the summation involves a pair of states: either a state $|u_{n}\rangle$ belonging to $P_{1}$ together with a state $|u_{\alpha}\rangle$ belonging to $Q$, or a state $|u_{n}\rangle$ belonging to $P$ together with a state $|u_{\alpha}\rangle$ belonging to $Q_{1}$. As illustrated schematically in Fig. S1, the two states in each pair are always separated by a continuously retained truncated subspace: $P_{n_{\rm cut}}$ in the former case, and $Q_{n_{\rm cut}}$ in the latter. As the cutoff $n_{\rm cut}$ is increased, both truncated subspaces expand, so that the denominator $E^{\rm HF}_{n}-E^{\rm HF}_{\alpha}$ correspondingly increases in magnitude. As a result, both differences $W_{\mu\nu}-W_{\mu\nu}^{\text{trun}}$ and $N_{\mu\nu}-N_{\mu\nu}^{\text{trun}}$ converge to zero. This guarantees the numerical stability of the present construction based on two truncated projectors.
It may be tempting instead to truncate only the occupied projector, denoted $P_{n_{\rm cut}}$, and to naively identify its complement with the projector onto the empty subspace, namely
$\widetilde{Q}_{n_{\rm cut}}(\bm{k})\equiv I-P_{n_{\rm cut}}(\bm{k})=Q_{n_{\rm cut}}(\bm{k})+Q_{1}(\bm{k})+P_{1}(\bm{k})\ .$ (S46)
Within this prescription, one immediately finds
$\widetilde{W}^{\rm trun}_{\mu\nu}(\bm{k})$ $=-\mathrm{Tr}\left[\big(\partial_{\mu}P_{n_{\rm cut}}\big)\big(Q_{n_{\rm cut}}+Q_{1}+P_{1}\big)\big(\partial_{\nu}P_{n_{\rm cut}}\big)\big(H^{\rm HF}-\mu\big)\right]$
$=W_{\mu\nu}^{\rm trun}+\mathrm{Tr}\left[\big(\partial_{\mu}P_{n_{\rm cut}}\big)Q_{1}\big(\partial_{\nu}P_{n_{\rm cut}}\big)\big(H^{\rm HF}-\mu\big)\right]+\mathrm{Tr}\left[\big(\partial_{\mu}P_{n_{\rm cut}}\big)P_{1}\big(\partial_{\nu}P_{n_{\rm cut}}\big)\big(H^{\rm HF}-\mu\big)\right]\ .$ (S47)
As discussed above, $W_{\mu\nu}^{\rm trun}$ is numerically stable because the occupied and empty subspaces are separated by the insulating gap. The second correction term,
$\mathrm{Tr}\left[\big(\partial_{\mu}P_{n_{\rm cut}}\big)Q_{1}\big(\partial_{\nu}P_{n_{\rm cut}}\big)\big(H^{\rm HF}-\mu\big)\right]$
is also stable, since the subspace $Q_{n_{\rm cut}}$ separates $P_{n_{\rm cut}}$ from $Q_{1}$, so that the energy interval between them increases as $n_{\rm cut}$ becomes larger. By contrast, no such interval exists between $P_{n_{\rm cut}}$ and $P_{1}$. More explicitly,
$\mathrm{Tr}\left[\big(\partial_{\mu}P_{n_{\rm cut}}\big)P_{1}\big(\partial_{\nu}P_{n_{\rm cut}}\big)\big(H^{\rm HF}-\mu\big)\right]$ $=\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}m>n_{\rm cut}\\ m\in\rm occ\end{subarray}}\langle\partial_{\mu}u_{n}|u_{m}\rangle\langle u_{m}|\partial_{\nu}u_{n}\rangle\left(E^{\rm HF}_{n}-\mu\right)$
$=\sum_{\begin{subarray}{c}n\leq n_{\rm cut}\\ n\in\rm occ\end{subarray}}\sum_{\begin{subarray}{c}m>n_{\rm cut}\\ m\in\rm occ\end{subarray}}\frac{\langle u_{n}|\partial_{\mu}H^{\rm HF}|u_{m}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{m}}\frac{\langle u_{m}|\partial_{\nu}H^{\rm HF}|u_{n}\rangle}{E^{\rm HF}_{n}-E^{\rm HF}_{m}}\left(E^{\rm HF}_{n}-\mu\right)\ .$ (S48)
As illustrated schematically in Fig. S1, the two subspaces $P_{n_{\rm cut}}$ and $P_{1}$ are directly connected, so that the denominator $E^{\rm HF}_{n}-E^{\rm HF}_{m}$ can, in principle, be arbitrarily small. Moreover, in the high-energy limit of a Dirac system, the density of states scales as $|E|$, so that the spacing between consecutive bands decreases approximately as $1/|E|$ as $n_{\rm cut}$ increases. As a result, the denominator $E^{\rm HF}_{n}-E^{\rm HF}_{m}$ becomes increasingly small, leading to large oscillatory contributions and consequently poor numerical convergence, as observed in scheme m1 shown in Fig. 2(b) of the main text.
The same issue also arises in approaches based on the covariant derivative, which is defined as Ceresoli et al. (2006)
$|\widetilde{\partial}_{\mu}u_{l}\rangle=\sum_{\alpha\neq l}|u_{\alpha}\rangle\langle u_{\alpha}|\partial_{\mu}u_{l}\rangle\ ,$ (S49)
if the complementary subspace is implemented through an incomplete projector of the form $I-P_{n_{\rm cut}}$. In that case, the projected derivative mixes the retained band with states lying on the same side of the gap, and the corresponding matrix elements are therefore no longer protected by the insulating gap.
Figure S1: Schematic illustration of the Hilbert-space decomposition responsible for the numerical stability of the symmetric truncation scheme. In panel (a), the retained occupied and empty subspaces, $P_{n_{\rm cut}}$ and $Q_{n_{\rm cut}}$, are separated by the insulating gap, so that the interband energy denominators entering $W_{\mu\nu}$ and $N_{\mu\nu}$ remain finite. In panel (b), the alternative choice $\widetilde{Q}_{n_{\rm cut}}=I-P_{n_{\rm cut}}$ additionally includes the omitted occupied sector $P_{1}$, therefore generating matrix elements across an artificial cutoff boundary within the occupied subspace. These terms are not gap-protected and therefore lead to unstable, oscillatory convergence.
V Derivation of the particle-hole symmetry of $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$
The same projector formalism also makes the particle-hole relation between the HF states at fillings $\nu$ and $\nu$ transparent. To establish the corresponding symmetry of the orbital magnetization, it is convenient to label the projectors and the self-consistent HF Hamiltonian by the filling, namely $P^{(\nu)}({\mathbf{k}})$, $Q^{(\nu)}({\mathbf{k}})$, and $H^{(\nu)}({\mathbf{k}})$. Let $\mathcal{P}$ denote the unitary particle-hole transformation of the BM model Bernevig et al. (2021). For a pair of particle-hole related HF solutions, $\mathcal{P}$ exchanges the occupied and empty subspaces according to
$\mathcal{P}P^{(\nu)}({\mathbf{k}})\mathcal{P}^{-1}$ $=Q^{(-\nu)}(-{\mathbf{k}}),$ (S50)
$\mathcal{P}Q^{(\nu)}({\mathbf{k}})\mathcal{P}^{-1}$ $=P^{(-\nu)}(-{\mathbf{k}}),$ (S51)
while the HF Hamiltonian, measured relative to the corresponding chemical potential, transforms as
$\mathcal{P}\big(H_{\nu}^{\mathrm{HF}}({\mathbf{k}})-\mu_{\nu}\big)\mathcal{P}^{-1}=-\big(H_{-\nu}^{\mathrm{HF}}(-{\mathbf{k}})-\mu_{-\nu}\big).$ (S52)
Substituting these relations into the operator definitions of $W_{\mu\nu}$ and $N_{\mu\nu}$ and using the invariance of the trace under a unitary transformation, we find
$W_{\mu\nu}^{(\nu)}({\mathbf{k}})$ $=\mathrm{Tr}\left[P^{(\nu)}({\mathbf{k}})\big(\partial_{\mu}P^{(\nu)}({\mathbf{k}})\big)\big(\partial_{\nu}Q^{(\nu)}({\mathbf{k}})\big)\big(H_{\nu}^{\mathrm{HF}}({\mathbf{k}})-\mu_{\nu}\big)\right]$
$=-\mathrm{Tr}\left[Q^{(-\nu)}(-{\mathbf{k}})\big(\partial_{\mu}Q^{(-\nu)}(-{\mathbf{k}})\big)\big(\partial_{\nu}P^{(-\nu)}(-{\mathbf{k}})\big)\big(H_{-\nu}^{\mathrm{HF}}(-{\mathbf{k}})-\mu_{-\nu}\big)\right]$
$=-\mathrm{Tr}\left[\big(\partial_{\mu}Q^{(-\nu)}(-{\mathbf{k}})\big)\big(\partial_{\nu}P^{(-\nu)}(-{\mathbf{k}})\big)Q^{(-\nu)}(-{\mathbf{k}})\big(H_{-\nu}^{\mathrm{HF}}(-{\mathbf{k}})-\mu_{-\nu}\big)\right]$
$=-\mathrm{Tr}\left[Q^{(-\nu)}(-{\mathbf{k}})\big(\partial_{\mu}P^{(-\nu)}(-{\mathbf{k}})\big)\big(\partial_{\nu}Q^{(-\nu)}(-{\mathbf{k}})\big)\big(H_{-\nu}^{\mathrm{HF}}(-{\mathbf{k}})-\mu_{-\nu}\big)\right]$
$=-N_{\mu\nu}^{(-\nu)}(-{\mathbf{k}})\ ,$ (S53)
and similarly
$N_{\mu\nu}^{(\nu)}({\mathbf{k}})=-W_{\mu\nu}^{(-\nu)}(-{\mathbf{k}})\ .$ (S54)
Substituting these identities into Eqs. 6 and 7 of the main text, and then summing over the mBZ, yields the general particle-hole relations
$M_{\mathrm{orb}}(\nu)=M_{\mathrm{orb}}(-\nu)\ ,\qquad m_{\mathrm{SR}}(\nu)=-m_{\mathrm{SR}}(-\nu)\ .$ (S55)
Importantly, the derivation above does not rely on the identity $P+Q=I$. The same symmetry relations therefore remain valid for the truncated quantities, provided that the truncation preserves particle-hole-related remote-band pairs symmetrically:
$M^{\rm trun}_{\mathrm{orb}}(\nu)=M^{\rm trun}_{\mathrm{orb}}(-\nu)\ ,\qquad m^{\rm trun}_{\mathrm{SR}}(\nu)=-m^{\rm trun}_{\mathrm{SR}}(-\nu)\ .$ (S56)
As shown in Fig. 2 and 3, the numerical results rapidly converge to the exact symmetry above as $n_{\mathrm{cut}}$ increases. In particular, for $\nu=\pm 3$ this gives the symmetry relations quoted in the main text and observed numerically in Fig. 3(c,d) of the main text.
VI Additional numerical results for benchmarking the formalism
In this section, we present additional numerical benchmarks for the projection-matrix formalism. We first repeat the HF and orbital-magnetization calculations for $\nu=-3$ on a denser $k$-point mesh with $N_{k}=60\times 60=3600$ points in the mBZ. The corresponding HF band structures are shown in Fig. S2(a,b), while the resulting $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ are displayed in Fig. S2(c,d). The results obtained from the two $k$-point meshes agree within 1% error, indicating that the calculation is well converged with respect to the $k$-mesh density in the mBZ.
Figure S2: (a,b) HF band structures of the Chern insulating state at filling $\nu=-3$, obtained using 900900 (a) and 36003600 (b) $k$ points in the mBZ. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). (c,d) Total orbital magnetization $M_{\mathrm{orb}}$ and self-rotation contribution $m_{\mathrm{SR}}$ per moiré cell for the states shown in (a) and (b), plotted as functions of the number of remote-band pairs $n_{\mathrm{cut}}$ included in the projection-matrix formalism. Both quantities are well converged with respect to the $k$-mesh density. When evaluating $M_{\mathrm{orb}}$, the chemical potential $\mu$ is fixed at the bottom of the conduction band, as indicated by the horizontal dashed lines in panels (a) and (b).
In the main text, for simplicity, we take the same cutoff $n_{\mathrm{cut}}$ for both projection operators $P_{n_{\rm cut}}(\mathbf{k})$ and $Q_{n_{\rm cut}}(\mathbf{k})$. More generally, the two cutoffs can be chosen independently. To illustrate this point, we repeat the calculation for $\nu=-3$ using $n_{{\rm cut},P}=n_{{\rm cut},Q}+10$ and $n_{{\rm cut},Q}=n_{{\rm cut},P}+10$. As shown in Fig. S3, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ exhibit convergence behavior very similar to the case with $n_{{\rm cut},P}=n_{{\rm cut},Q}$.
Figure S3: Total orbital magnetization $M_{\mathrm{orb}}$ (a) and self-rotation contribution $m_{\mathrm{SR}}$ (b) per moiré cell for the Chern insulating state at $\nu=-3$, plotted as functions of the independently chosen remote-band cutoffs $n_{{\rm cut},P}$ and $n_{{\rm cut},Q}$ in the projection-matrix formalism. Both quantities show clear convergence around $n_{\mathrm{cut}}\approx 20$, even when $n_{{\rm cut},P}\neq n_{{\rm cut},Q}$.
VII Orbital magnetization for $\nu=\pm 2$
We next present HF results for fillings $\nu=\pm 2$. Owing to the particle-hole symmetry relating $\nu$ and $\nu$, it is sufficient to focus on $\nu=2$. The Hartree–Fock calculation yields two nearly degenerate correlated insulating states: a correlated Chern insulator (CCI) and an intervalley-coherent (IVC) insulator. The lower-energy solution is the IVC state, whose bands are topologically trivial, as shown in Fig. S4(b). It is lower in energy by approximately 0.4 meV per moiré cell than the CCI state, in which two valence bands break $C_{2}\mathcal{T}$ symmetry and each carry a finite Chern number. Among the degenerate CCI solutions, we show the state with total Chern number $C=2$, whose band dispersion is plotted in Fig. S4(a).
We then evaluate the orbital magnetization for both states. As shown in Fig. S4(c), the total orbital magnetization $M_{\mathrm{orb}}$ of the CCI depends linearly on the chemical potential within the insulating gap and changes sign across the gap, whereas $M_{\mathrm{orb}}$ vanishes for the IVC state because its bands are topologically trivial. This indicates that the CCI is energetically favored over the IVC state in the presence of a finite magnetic field. The self-rotation contribution $m_{\mathrm{SR}}$, shown in Fig. S4(d), is likewise sizable for the CCI but vanishes for the IVC state. Using the magnetization of CCI in the center of the gap and the energy difference between the two insulating states, we also estimate a critical magnetic field for the transition from the IVC state to the CCI of $B_{c}\sim\delta E/\delta M_{\mathrm{orb}}\approx 9.2$ T.
Figure S4: (a,b) HF band dispersions of the correlated Chern insulating state (CCI) (a) and the intervalley-coherent (IVC) state (b) at filling $\nu=2$. The solid curves represent the eight interaction-renormalized HF bands, while the dashed curves denote the remote bands excluded from the self-consistent HF calculation. Bands carrying nontrivial Chern numbers are highlighted in red ($C=1$) and blue (C=1C=-1). For the CCI, the Chern bands are doubly degenerate, yielding a total Chern number $C=2$. (c,d) Total orbital magnetization $M_{\mathrm{orb}}$ (c) and self-rotation contribution $m_{\mathrm{SR}}$ (d) per moiré cell as functions of the chemical potential $\mu$ inside the insulating gap for the two states shown in (a) and (b). Solid circles correspond to the CCI, and solid squares to the IVC state. For the CCI, $M_{\mathrm{orb}}$ varies linearly with $\mu$ and changes sign across the gap, while $m_{\mathrm{SR}}$ remains constant. For the IVC state, both $M_{\mathrm{orb}}$ and $m_{\mathrm{SR}}$ vanish because all bands are topologically trivial. The number of remote-band pairs is fixed at $n_{\mathrm{cut}}=30$ in the calculation.